### v2

# V. Mehrmann and E. MengiRobust Stability of a Quadratic Matrix Polynomial

Abstract
For a quadratic matrix polynomial dependent on parameters and a given tolerance $\epsilon>0$, the minimization of the $\epsilon$-pseudospectral abscissa over the set of permissible parameter values is discussed, with applications in damping optimization and brake squeal reductions in mind. An approach is introduced that is based on nonsmooth and global optimization (or smooth optimization techniques such as BFGS if there are many parameters) equipped with a globally convergent criss-cross algorithm to compute the $\epsilon$-pseudospectral abscissa objective when the matrix polynomial is of small size. For the setting when the matrix polynomial is large, a subspace framework is introduced, and it is argued formally that it solves the minimization problem globally. The subspace framework restricts the parameter-dependent matrix polynomial to small subspaces, and thus solves the minimization problem for such restricted small matrix polynomials. It then expands the subspaces using the minimizers for the restricted polynomials. The proposed approach makes the global minimization of the $\epsilon$-pseudospectral abscissa possible for a quadratic matrix polynomial dependent on a few parameters and for sizes up to at least a few hundreds. This is illustrated on several examples originating from damping optimization.
keywords:
quadratic eigenvalue problem, pseudospectral abscissa, damping optimization, robust stability, subspace framework, nonsmooth optimization, global optimization
{AMS}
65F15, 93D09, 90C26
1 Introduction
We consider a quadratic matrix polynomial that depends on parameters, and, in particular, we are interested in minimizing the $\epsilon$-pseudospectral abscissa of such a matrix polynomial over the parameters. Our work is motivated by applications such as damping optimization [29, 34] and the brake squealing problem [10]. Having a negative $\epsilon$-pseudospectral abscissa after minimization guarantees robust stability, i.e. the asymptotic stability of not only the associated second-order linear time-invariant system, but also all nearby systems.
The parameter-dependent matrix polynomial under consideration is of the form
(1) $P(\lambda;\nu)\;=\;\lambda^{2}M(\nu)\>+\>\lambda C(\nu)\>+\>K(\nu)$
where $\nu\in\Omega\subseteq{\mathbb{R}}^{\mathsf{d}}$ represents a vector of parameters, $\Omega$ is an open subset of ${\mathbb{R}}^{\mathsf{d}}$, and the coefficients have the special form
(2) $\begin{split}&M(\nu)\;=\;\>\psi_{1}(\nu)M_{1}+\dots+\psi_{\mathsf{m}}(\nu)M_{\mathsf{m}}\,,\\[3.00003pt] &C(\nu)\;\>=\;\>\zeta_{1}(\nu)C_{1}+\dots+\zeta_{\mathsf{c}}(\nu)C_{\mathsf{c}}\,,\\[3.00003pt] &K(\nu)\;\,=\;\>\kappa_{1}(\nu)K_{1}+\dots+\kappa_{\mathsf{k}}(\nu)K_{\mathsf{k}}\,,\end{split}$
for given (parameter-independent) matrices $M_{i},C_{j},K_{\ell}\in{\mathbb{C}}^{\mathsf{n}\times\mathsf{n}}$, and real-analytic functions $\psi_{i},\zeta_{j},\kappa_{\ell}:\Omega\rightarrow{\mathbb{R}}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\>\ell=1,\dots,\mathsf{k}$. In a context where the dependence of $P(\lambda;\nu)$ on $\nu$ is not important, we may leave out this explicit dependence on $\nu$, and simply write
(3) $P(\lambda)\;=\;\lambda^{2}M+\lambda C+K$
for matrices $M,C,K\in{\mathbb{C}}^{\mathsf{n}\times\mathsf{n}}$.
There are slightly different (but closely related) definitions of the $\epsilon$-pseudospectrum of the matrix polynomial $P(\lambda)$ in (3) in the literature, [22, 32, 33]. In this work, for computational convenience, we use the definition from [22]:
(4) $\begin{split}\Lambda_{\epsilon}(P)\;:=&\;\{\lambda\in{\mathbb{C}}\;|\;\text{det}(P(\lambda)+\Delta P(\lambda))=0\\[1.99997pt] &\;\;\;\;\;\text{ for some }\Delta P(\lambda)=w_{m}\lambda^{2}\Delta M+w_{c}\lambda\Delta C+w_{k}\Delta K\\[1.99997pt] &\;\;\;\;\;\text{ with }\,\Delta M,\Delta C,\Delta K\in{\mathbb{C}}^{\mathsf{n}\times\mathsf{n}}\;\text{ s.t. }\|\left[\begin{array}[]{ccc}\Delta M&\Delta C&\Delta K\end{array}\right]\|_{2}\leq\epsilon\}\;,\end{split}$
with $\|\cdot\|_{2}$ representing the matrix 2-norm, and nonnegative real scalars $w_{m},w_{c},w_{k}$ that can be regarded as weights. It follows from the arguments in the proof of [8, Lemma 8] that $\Lambda_{\epsilon}(P)$ can be characterized as
(5) $\Lambda_{\epsilon}(P)\;=\;\{\lambda\in{\mathbb{C}}\;|\;\sigma_{\min}(P(\lambda))\leq\epsilon p_{w}(|\lambda|)\}$
where $\sigma_{\min}(A)$ denotes the smallest singular value of the matrix $A$, and
(6) $p_{w}(z)\;:=\;\sqrt{w_{m}^{2}z^{4}+w_{c}^{2}z^{2}+w_{k}^{2}},$
as well as $w:=(w_{m},w_{c},w_{k})$.
Extended definitions of the $\epsilon$-pseudospectrum for more general polynomial eigenvalue problems, as well as general nonlinear eigenvalue problems are also presented in the literature [22, 26, 32, 33].
The $\epsilon$-pseudospectral abscissa of $P$ is defined as
(7) $\begin{split}\alpha_{\epsilon}(P)\;&:=\;\sup\{\mathrm{Re}(z)\;|\;z\in\Lambda_{\epsilon}(P)\}\\[3.00003pt] \;&\phantom{:}=\;\sup\{\mathrm{Re}(z)\;|\;z\in{\mathbb{C}}\text{ s.t. }\sigma_{\min}(P(z))\leq\epsilon p_{w}(|z|)\}.\end{split}$
The main problem that we deal with is the minimization of the $\epsilon$-pseudospectral abscissa of $P(\lambda;\nu)$ in (1) over the set of permissible parameter values, i.e.
$\min\{\,\alpha_{\epsilon}(P(\,\cdot\,;\nu))\;|\;\nu\in\underline{\Omega}\,\}$
with $\underline{\Omega}\subseteq\Omega$ denoting a compact set of permissible parameter values. We propose approaches for both small-scale problems when $\mathsf{n}$ is small, and moderate- to large-scale problems when $\mathsf{n}$ is not small. Throughout this work, we keep the following boundedness assumption.
[…]
$\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ is bounded for all $\nu\in\underline{\Omega}$.
Under this boundedness assumption, a rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ must exist, i.e. the supremum in the definition of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ (see (7) above) must be attained, so it can be replaced by the maximum for all $\nu\in\underline{\Omega}$.
The ideas and approaches presented in this paper can be extended in a straightforward way to a parameter-dependent nonlinear eigenvalue problem when $P(\lambda;\nu)$ in (1) is replaced by a more general matrix-valued function
(8) $T(\lambda;\nu)\;=\;g_{0}(\lambda)A_{0}(\nu)\>+\>g_{1}(\lambda)A_{1}(\nu)\>+\>\dots\>+\>g_{\ell}(\lambda)A_{\ell}(\nu)$
for given meromorphic functions $g_{j}:{\mathcal{D}}\rightarrow{\mathbb{C}}$ for $j=0,1,\dots,\ell$ with a domain ${\mathcal{D}}$ that is a dense subset of ${\mathbb{C}}$. The polynomial eigenvalue problem is a special case of (8) with $g_{j}(\lambda)=\lambda^{j}$ for $j=0,1,\dots,\ell$.
To have a concise presentation, we restrict ourselves to quadratic matrix polynomials with $g_{j}(\lambda)=\lambda^{j}$ for $j=0,1,2$. The outcome of our work is an efficient approach to optimize the robust stability of a parameter-dependent quadratic eigenvalue problem, viewed as an important challenging problem for instance by the damping optimization community [27, 28, 34]. The new approach is designed systematically, especially well-suited for large problems, and has theoretical global convergence guarantee.
The current work can be regarded as a generalization of [2] that concerns the minimization of the $\epsilon$-pseudospectral abscissa of a matrix dependent on parameters. We extend the analytical properties of the $\epsilon$-pseudospectral abscissa functions from the matrix setting to the quadratic matrix polynomial setting. The algorithms are built on top of these analytical properties, and implementing them requires considerable effort and involves many subtleties.
Our method comes with publicly available MATLAB implementations [21] of the proposed algorithms. We illustrate the effectiveness of the proposed algorithms in practice by applying them to real damping optimization problems.
Outline. In Section 2 we propose algorithms for small problems involving coefficients of small dimension $\mathsf{n}$. We first describe a globally convergent criss-cross algorithm to compute the $\epsilon$-pseudospectral abscissa and the rightmost point in the $\epsilon$-pseudospectrum. Then, we discuss how $\alpha_{\epsilon}(P(\cdot;\nu))$ can be minimized in the small-scale setting, where we present conditions ensuring the real analyticity of $\alpha_{\epsilon}(P(\cdot;\nu))$, and derive expressions for the derivatives of $\alpha_{\epsilon}(P(\cdot;\nu))$.
Section 3 is devoted to large-scale computation of the $\epsilon$-pseudospectral abscissa and the corresponding rightmost point, in particular to the development of a subspace framework for this purpose when the dimension of the matrix polynomial is not small.
Section 4 focuses on the minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$ in the case that the parameter-dependent matrix polynomial at hand is not of small dimension. A second subspace framework for minimizing $\alpha_{\epsilon}(P(\cdot;\nu))$ is introduced, and a formal argument is given in support of its convergence to a global minimizer of $\alpha_{\epsilon}(P(\cdot;\nu))$.
Section 5 concerns the applications of the proposed approaches for $\epsilon$-pseudospectral abscissa minimization to several examples originating from damping optimization. An alternative to directly minimizing $\alpha_{\epsilon}(P(\cdot;\nu))$ over $\nu$ is minimizing the $\epsilon$-pseudospectral abscissa of a linearization of $P(\cdot;\nu)$ over $\nu$. We provide a comparison of these two $\epsilon$-pseudospectral abscissa minimization problems in Section 6, and an example for which the minimizers of these two minimization problems differ considerably.
We summarize our results, and potential directions for future research in Section 7.
2 Small-scale problems
In this section, we deal with a matrix polynomial that has small size, and describe approaches for the computation and minimization of its $\epsilon$-pseudospectral abscissa.
2.1 Computation of the $\epsilon$-pseudospectral abscissa for small-scale problems
Inspired by the ideas in [5] for the matrix case, in [22] a criss-cross algorithm is presented to compute the $\epsilon$-pseudospectral abscissa $\alpha_{\epsilon}(P)$ for a small-scale matrix polynomial $P(\lambda)$. We briefly recall this algorithm starting with its two main ingredients, namely a vertical search and a horizontal search.
The vertical search for a given $x\in{\mathbb{R}}$, stated below as a theorem, is helpful to locate the intersection points of the boundary of $\Lambda_{\epsilon}(P)$ in the complex plane with the vertical line $\mathrm{Re}(z)=x$.
Theorem 2.1 (Vertical Search [22]).
For every $x\in{\mathbb{R}}$, the following statements are equivalent:
  1. (i)
    One of the singular values of $\,\{1/p_{w}(|x+{\rm i}y|)\}P(x+{\rm i}y)$ for $y\in{\mathbb{R}}$ equals $\epsilon$.
  2. (ii)
    The matrix polynomial ${\mathcal{M}}(\lambda;x)=\sum_{j=0}^{4}\lambda^{j}{\mathcal{M}}_{j}(x)$ with
    ${\mathcal{M}}_{j}(x)\;:=\;\left\{\begin{array}[]{cl}\left[\begin{array}[]{cc}-\epsilon p^{2}_{w}(x)I&\;[P(x)]^{\ast}\\[2.5pt] P(x)&-\epsilon I\end{array}\right]&\quad\text{if }j=0\;,\\[13.00005pt] \left[\begin{array}[]{cc}0&\;-[P^{\prime}(x)]^{\ast}\\[2.5pt] P^{\prime}(x)&\;0\end{array}\right]&\quad\text{if }j=1\;,\\[16.00008pt] \left[\begin{array}[]{cc}\epsilon(2w_{m}^{2}x^{2}+w_{c}^{2})I&\;M^{\ast}\\[2.5pt] M&0\end{array}\right]&\quad\text{if }j=2\;,\\[18.99995pt] 0&\quad\text{if }j=3\;,\\[13.00005pt] \left[\begin{array}[]{cc}-\epsilon w_{m}^{2}I&\;0\\ 0&\;0\end{array}\right]&\quad\text{if }j=4\end{array}\right.$
    has ${\rm i}y$ as an eigenvalue.
According to Theorem 2.1, the set of intersection points of the boundary of $\Lambda_{\epsilon}(P)$ with the vertical line $\mathrm{Re}(z)=x$ can be obtained by extracting the purely imaginary eigenvalues ${\rm i}y_{1},\dots\rm iy_{\eta(x)}$ of ${\mathcal{M}}(\lambda;x)$, then checking whether
$\sigma_{\min}(P(x+{\rm i}y_{j}))\;=\;\epsilon\left\{p_{w}(|x+{\rm i}y_{j}|)\right\}$
for $j=1,\dots,\eta(x)$.
We next state, as a theorem, the horizontal search, which concerns finding the rightmost intersection point of the $\epsilon$-pseudospectrum with a horizontal line in the complex plane.
Theorem 2.2 (Horizontal Search [22]).
For a given $y\in{\mathbb{R}}$, if the set
${\mathcal{H}}(y)\;:=\;\Lambda_{\epsilon}(P)\cap\left\{\mathrm{Im}(z)=y\right\}$
is not empty, then the matrix polynomial ${\mathcal{N}}(\lambda;y)=\sum_{j=0}^{4}\lambda^{j}{\mathcal{N}}_{j}(y)$ with
${\mathcal{N}}_{j}(y)\;:=\;\left\{\begin{array}[]{cl}\left[\begin{array}[]{cc}-\epsilon p^{2}_{w}(y)I&\;\;[P({\rm i}y)]^{\ast}\\[2.5pt] P({\rm i}y)&-\epsilon I\end{array}\right]&\quad\text{if }j=0\;,\\[13.00005pt] \left[\begin{array}[]{cc}0&\;-{\rm i}[P^{\prime}({\rm i}y)]^{\ast}\\[2.5pt] -{\rm i}P^{\prime}({\rm i}y)&\;\;0\end{array}\right]&\quad\text{if }j=1\;,\\[16.00008pt] \left[\begin{array}[]{cc}\epsilon(2w_{m}^{2}y^{2}+w_{c}^{2})I&\;-M^{\ast}\\[2.5pt] -M&\;\;0\end{array}\right]&\quad\text{if }j=2\;,\\[18.99995pt] 0&\quad\text{if }j=3\;,\\[13.00005pt] \left[\begin{array}[]{cc}-\epsilon w_{m}^{2}I&\;0\\ 0&\;0\end{array}\right]&\quad\text{if }j=4\end{array}\right.$
has a purely imaginary eigenvalue. Moreover, if ${\mathcal{H}}(y)$ is not empty, then
(9) $\max\,\mathrm{Re}\{{\mathcal{H}}(y)\}\;=\;\max\{x\;|\;x\in{\mathbb{R}}\,\text{s.t. }\,{\rm i}x\in\Lambda({\mathcal{N}}(\lambda;y))\}$
where $\Lambda({\mathcal{N}}(\lambda;y))$ denotes the set of finite eigenvalues of the matrix polynomial ${\mathcal{N}}(\lambda;y)$.
According to Theorem 2.2, the rightmost intersection point of $\Lambda_{\epsilon}(P)$ with the horizontal line $\mathrm{Im}(z)=y$ in the complex plane can be determined from the set of purely imaginary eigenvalues of ${\mathcal{N}}(\lambda;y)$ by employing (9).
The criss-cross algorithm to compute $\alpha_{\epsilon}(P)$ at every iteration for a given estimate $\underline{x}$ first finds the intervals on the line $\mathrm{Re}(z)=\underline{x}$ that lie in $\Lambda_{\epsilon}(P)$ by means of a vertical search. Then a horizontal search is performed at the midpoint of each such interval, and the estimate $\underline{x}$ is updated to the real part of the rightmost point in $\Lambda_{\epsilon}(P)$ obtained from these horizontal searches. A formal description of the approach is given in Algorithm 1.
0:  A quadratic matrix polynomial $P(\lambda)$ of the form (3), and $\epsilon>0$.
0:  Estimates $\underline{x}$ for $\alpha_{\epsilon}(P)$, and $\underline{z}\in{\mathbb{C}}$ for a rightmost point in $\Lambda_{\epsilon}(P)$.
1:  $\underline{z}\leftarrow$ a rightmost eigenvalue of $P(\lambda)$.
2:  $\underline{x}^{(0)}\leftarrow$ $\text{Re}(\underline{z})$.
3:  for $k=1,2,\dots$ do
4:     Vertical Scan. Determine the intervals $[y^{L}_{j},y^{R}_{j}]$ for $j=1,\dots,l$ such that
(10) $\bigcup_{j=1}^{l}[y^{L}_{j},y^{R}_{j}]\;=\;\mathrm{Im}\left\{\Lambda_{\epsilon}(P)\cap\left\{\mathrm{Re}(z)=\underline{x}^{(k-1)}\right\}\right\}$
by means of a vertical search at $\underline{x}^{(k-1)}$.
5:     Horizontal Refinement. Letting $y^{\mathrm{mid}}_{j}:=(y^{L}_{j}+y^{R}_{j})/2$, set
(11) $\underline{x}^{(k)}\leftarrow\max_{j=1,\dots,l}\max\left\{\mathrm{Re}\left\{\Lambda_{\epsilon}(P)\cap\left\{\mathrm{Im}(z)=y^{\mathrm{mid}}_{j}\right\}\right\}\right\}$
by means of horizontal searches at $y^{\mathrm{mid}}_{j}$ for $j=1,\dots,l$, and set $\,\underline{z}\leftarrow\underline{x}^{(k)}+{\rm i}\,y^{\mathrm{mid}}_{\underline{j}}\,$ with $\underline{j}$ denoting the maximizing $j$ in (11).
6:     Return $\underline{x}\leftarrow\text{Re}(\underline{z})\>$ and $\,\underline{z}\;$ if convergence occurred.
7:  end for
Algorithm 1 The criss-cross algorithm to compute $\alpha_{\epsilon}(P)$
An example of the progress of Algorithm 1 is presented for a $4\times 4$ quadratic matrix polynomial in Figure 1. In this example, the algorithm first finds the intersection points of the line $\mathrm{Re}(z)=\alpha(P)$ with the boundary of $\Lambda_{\epsilon}(P)$ by means of a vertical search, where $\alpha(P)$ denotes the spectral abscissa of $P$, i.e. the real part of the rightmost eigenvalue of $P$ (see the left-hand plot). Using these intersection points, it determines the three intervals on the line $\mathrm{Re}(z)=\alpha(P)$ that lie inside $\Lambda_{\epsilon}(P)$, and performs a horizontal search at the midpoint of each one of these three intervals (see the middle plot). Then the algorithm performs another vertical search to find the intersection points of the boundary of $\Lambda_{\epsilon}(P)$ with the line $\mathrm{Re}(z)=\underline{x}$, where $\underline{x}$ is the real part of the rightmost point returned by the three horizontal searches (see the right-hand plot).
Refer to caption Refer to caption Refer to caption
Figure 1: Progress of Algorithm 1 for a $4\times 4$ quadratic matrix polynomial. The blue curves correspond to the boundary $\partial\Lambda_{\epsilon}$ of the $\epsilon$-pseudospectrum of the quadratic matrix polynomial for $\epsilon=0.7$.
Convergence. It has been shown in [5] that the estimate $x^{(k)}$ generated by Algorithm 1 in the matrix case converges to the $\epsilon$-pseudospectral abscissa as $k\rightarrow\infty$, and the rate of this convergence is at least quadratic. These results extend to Algorithm 1, that is to the criss-cross algorithm for the more general quadratic matrix polynomial setting. The proofs of the convergence results stated below for Algorithm 1 are similar to their counterparts in the matrix setting, in particular to [5, Theorem 3.2] and [5, Theorem 5.2], but by replacing the singular value function $h(x,y)$ in there with
(12) $g(x,y)\;:=\;\sigma_{\min}(P(x+{\rm i}y))-\epsilon p_{w}(\sqrt{x^{2}+y^{2}}).$
These results are stated assuming there is no termination, i.e. for Algorithm 1 without line 6.
Theorem 2.3 (Global Convergence).
The sequence $\{x^{(k)}\}$ generated by Algorithm 1 is monotonically non-decreasing and satisfies $\>\lim_{k\rightarrow\infty}x^{(k)}=\alpha_{\epsilon}(P)$.
Theorem 2.4 (At Least Quadratic Convergence).
Suppose that at every globally rightmost point $z_{\ast}=x_{\ast}+{\rm i}y_{\ast}$ with $x_{\ast},y_{\ast}\in{\mathbb{R}}$ of $\Lambda_{\epsilon}(P)$ the singular value $\sigma_{\min}(P(z_{\ast}))$ is simple and $\nabla g(x_{\ast},y_{\ast})\neq 0$. Then the sequence $\{x^{(k)}\}$ generated by Algorithm 1 satisfies
$|x^{(k+1)}-\alpha_{\epsilon}(P)|\;\leq\;{\mathcal{C}}|x^{(k)}-\alpha_{\epsilon}(P)|^{2}$
for some constant ${\mathcal{C}}\geq 0$.
Table 1 lists the iterates of the algorithm on two exemplary quadratic matrix polynomials. For both polynomials the iterates converge at a quadratic rate as expected.
$k$ $x^{(k)}$
0 0.152872406397
1 0.422554211904
2 0.426672961073
3 0.426804683720
4 0.426804822541
$k$ $x^{(k)}$
0 0.731311985577
1 0.855240708191
2 0.855256225558
3 0.855256225560
Table 1: Lists of iterates of Algorithm 1 on two quadratic matrix polynomials. Left: Polynomial of size $4\times 4$, $\epsilon=0.7$ and $(w_{m},w_{c},w_{k})=(1,1,1)$. Right: Polynomial of size $100\times 100$, $\epsilon=0.3$ and $(w_{m},w_{c},w_{k})=(0.4,0.3,1)$.
Termination. We terminate the algorithm in line 6 if the estimate $x^{(k)}$ after update in line 5 is not greater than the previous estimate $x^{(k-1)}$. Another possibility for termination is that the vertical search in line 4 cannot find any intersection points of $\Lambda_{\epsilon}(P)$ with $\mathrm{Re}(z)=x^{(k-1)}$, i.e. the matrix polynomial ${\mathcal{M}}(\lambda;x^{(k-1)})$ does not have any imaginary eigenvalue. In this second possibility, we simply skip the horizontal refinement in line 5 and terminate in line 6.
Computational cost. The computationally most expensive parts of Algorithm 1 are the vertical search in line 4 and $l$ horizontal searches in line 5 that need to be performed at every iteration. Each vertical and horizontal search typically requires the computation of all eigenvalues of a quartic matrix polynomial of size $2\mathsf{n}\times 2\mathsf{n}$, that can be determined at an ${\mathcal{O}}(\mathsf{n}^{3})$ computational cost in practice. Hence, every iteration of Algorithm 1 has a cubic cost ${\mathcal{O}}(\mathsf{n}^{3})$ in practice, yet the constant hidden in the ${\mathcal{O}}$-notation is large, since the eigenvalue problems are of size $2\mathsf{n}\times 2\mathsf{n}$ and involve quartic matrix polynomials. This makes the algorithm suitable only for small matrix polynomials.
2.2 Minimization of the $\epsilon$-pseudospectral abscissa for small-scale problems
The pseudospectral abscissa function $\alpha_{\epsilon}(\nu):=\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ is usually not convex in the parameters $\nu$. Moreover, it is typically not smooth whenever the rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ is not unique, or the singular value $\sigma_{\min}(P(\underline{z}\,;\nu))$ at the rightmost point $\underline{z}$ of $\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ is not simple. Smooth optimization techniques, especially BFGS, have been used successfully on such nonsmooth optimization problems; see in particular [6] and the MATLAB package GRANSO, which is based on BFGS. Such smooth optimization techniques operate locally, and can at best be expected to converge to a local minimizer of $\alpha_{\epsilon}(\nu)$, that is likely not a global minimizer. On the other end, global optimization techniques may guarantee convergence to a global minimizer of $\alpha_{\epsilon}(\nu)$, but they become computationally infeasible if there are more than a few parameters. Such a global optimization technique based on global piecewise quadratic approximations [4] is adopted for eigenvalue functions in [25]. Its implementation EigOpt is publicly available. At every iteration, this approach minimizes the maximum of several quadratic functions bounding the objective from below and interpolating the objective at several points, then adds a new quadratic function that interpolates the objective at the computed minimizer to existing quadratic functions, and again minimizes the maximum of quadratic functions. We emphasize that both GRANSO and EigOpt work well for nonsmooth problems, as long as the objective is continuous, piecewise continuously differentiable and differentiable almost everywhere, which is the case for $\alpha_{\epsilon}(\nu)$. In particular, they are both well-suited to converge to nonsmooth minimizers of $\alpha_{\epsilon}(\nu)$ provided that $\alpha_{\epsilon}(\nu)$ is differentiable at the iterates of these algorithms. We refer to [17] and [25, Theorem 8.1] for the convergence theory of GRANSO and EigOpt, respectively, when the objective is nonsmooth at the minimizer.
In the small-scale setting, we minimize $\alpha_{\epsilon}(\nu)$ either with GRANSO if there are many parameters, or with EigOpt if there are only a few parameters. Note that EigOpt at termination guarantees that the $\epsilon$-pseudospectral abscissa objective at the computed estimate for the minimizer does not differ from the actual globally smallest value of the objective by more than a prescribed tolerance. On the other hand, for GRANSO there are several termination criteria, but in our context here it normally terminates either when the norm of the search direction is less than a prescribed tolerance, or the line-search in the search direction fails to find a point satisfying the weak Wolfe conditions (i.e. that consist of a sufficient decrease and a sufficient curvature condition [30]). Both methods require not only the computation of $\alpha_{\epsilon}(\nu)$, but also the first derivatives of $\alpha_{\epsilon}(\nu)$. Fortunately, once $\alpha_{\epsilon}(\nu)$ as well as the rightmost point where it is attained are computed by Algorithm 1, then the first derivatives of $\alpha_{\epsilon}(\nu)$ can be obtained at almost no additional computational cost.
To see this, for $P(\nu,s):=P(s_{1}+{\rm i}s_{2};\nu)$ and $\sigma(\nu,s):=\sigma_{\min}(P(\nu,s))$ with $s=(s_{1},s_{2})\in{\mathbb{R}}^{2}$, the $\epsilon$-pseudospectral abscissa function under consideration can be expressed as
(13) $\alpha_{\epsilon}(\nu)\;=\;\max\{s_{1}\;|\;s=(s_{1},s_{2})\in{\mathbb{R}}^{2}\;\text{ is s.t. }\;\sigma(\nu,s)-\epsilon p_{w}(\|s\|)\leq 0\}$
with $\|s\|:=\sqrt{s_{1}^{2}+s_{2}^{2}}$ denoting the Euclidean norm of $s$, and an associated Lagrangian function (in the primal variable $s$ and the dual variable $\mu$) is given by
${\mathcal{L}}(\nu,s,\mu)\;:=\;s_{1}-\mu\{\sigma(\nu,s)-\epsilon p_{w}(\|s\|)\}\;.$
For simplicity, we combine the primal and dual variable into a vector $y=(s,\mu)$. Moreover, we denote by ${\mathcal{L}}_{s_{1}}$, ${\mathcal{L}}_{s_{2}}$ and ${\mathcal{L}}_{y}$, ${\mathcal{L}}_{yy}$ the partial derivative of ${\mathcal{L}}$ with respect to $s_{1}$, $s_{2}$ and the gradient, the Hessian of ${\mathcal{L}}$ with respect to $y$. If the optimizer $s(\nu)$ of (13) is unique and $\sigma(\nu,s(\nu))$ is simple, then, assuming the satisfaction of a constraint qualification, there exists a unique $\mu(\nu)$ such that ${\mathcal{L}}_{y}(\nu,s(\nu),\mu(\nu))=0$; see the proof of Theorem 2.6 below. We also combine the optimal values of the primal and dual variables, and write $y(\nu)=(s(\nu),\mu(\nu))$.
The subsequent derivations make use of the smoothness result concerning the singular value functions given below. This can be regarded as a special case of the classical results related to the eigenvalue functions of Hermitian matrix-valued functions that depend on its parameters in a real-analytic manner [13, 15, 31]. The extension of these results to singular value functions has been discussed in [25, Section 3.3].
Lemma 2.5.
Let $A:{\mathbb{R}}^{\ell}\rightarrow{\mathbb{C}}^{p\times q}$ be a matrix-valued function such that its real part $(1/2)\{A+\overline{A}\}$ as well as imaginary part $(-{\rm i}/2)\{A-\overline{A}\}$ are real-analytic, and the $j$th largest singular value $\sigma_{j}(\mu)$ of $A(\mu)$ is simple and nonzero at $\mu=\widetilde{\mu}$. Then $\sigma_{j}(\mu)$ is real-analytic at $\mu=\widetilde{\mu}$, and its first derivatives are given by
(14) $\frac{\partial\sigma_{j}}{\partial\mu_{k}}(\widetilde{\mu})\;=\;{\mathrm{R}e}\left\{u^{\ast}_{j}\,\frac{\partial A}{\partial\mu_{k}}(\widetilde{\mu})\,v_{j}\right\}$
for $k=1,\dots,\ell$, where $u_{j},v_{j}$ denote consistent unit left, respectively right singular vectors of $A(\widetilde{\mu})$ corresponding to the singular value $\sigma_{j}(\widetilde{\mu})$.
We refer to a point $\nu\in\Omega$ as a non-degenerate point of (13) if
  1. 1.
    there is a unique optimizer of (13), denoted as $s(\nu)=(s_{1}(\nu),s_{2}(\nu))$,
  2. 2.
    the singular value $\sigma(\nu,s(\nu))$ of $P(\nu,s(\nu))$ is simple, and
  3. 3.
    the Hessian ${\mathcal{L}}_{yy}(\nu,y(\nu))$ is invertible.
As we will prove in the next result, the three conditions above combined with a mild assumption guarantee the smoothness of $\alpha_{\epsilon}(\nu)$. For the subsequent result, $P^{\prime}(s;\nu)$ stands for the ordinary derivative of the polynomial $P(\,\cdot\,;\nu)$ given by $P^{\prime}(s;\nu)=2sM(\nu)+C(\nu)$. Additionally, $g$ stands for the function defined in (12), but to make its dependence on the parameters $\nu$ explicit, we use
(15) $g(s;\nu)\;:=\;\sigma(\nu,s)-\epsilon p_{w}(\|s\|).$
Also, $g_{s_{1}}$, $g_{s_{2}}$ represent the partial derivatives of $g$ with respect to $s_{1}$, $s_{2}$, and $g_{s}$ the gradient of $g$ with respect to $s$.
Theorem 2.6.
Let $\widetilde{\nu}$ be a non-degenerate point of (13). Moreover, suppose that $p_{w}(\|s(\widetilde{\nu})\|)\neq 0\,$ and $\,g_{s}(s(\widetilde{\nu});\widetilde{\nu})\neq 0\,$. Then the function $\nu\mapsto\alpha_{\epsilon}(\nu)$ is real analytic at $\widetilde{\nu}$ with the first derivatives given by
(16) $\frac{\partial\alpha_{\epsilon}}{\partial\nu_{j}}(\widetilde{\nu})\;=\;-\mu(\widetilde{\nu})\mathrm{Re}\left\{u^{\ast}P_{\nu_{j}}(\widetilde{\nu},s(\widetilde{\nu}))v\right\}\,,\;\;\;\;j=1,\dots,\mathsf{d}\,,$
where $P_{\nu_{j}}$ denotes the partial derivative of $P$ with respect to $\nu_{j}$,
(17) $\mu(\widetilde{\nu})\;:=\;\frac{1}{\mathrm{Re}\{u^{\ast}P^{\prime}(s(\widetilde{\nu});\widetilde{\nu})v\}-\epsilon\{s_{1}(\widetilde{\nu})/\|s(\widetilde{\nu})\|\}p^{\prime}_{w}(\|s(\widetilde{\nu})\|)}\,,$
and $u,v$ are consistent unit left, right singular vectors, respectively, of $P(\widetilde{\nu},s(\widetilde{\nu}))$ corresponding to its smallest singular value $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$.
Proof 2.7.
As $\widetilde{\nu}$ is non-degenerate, (13) with $\nu=\widetilde{\nu}$ has a unique optimizer, which we denote by $s(\widetilde{\nu})$, and $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$ is simple. If $\sigma(\widetilde{\nu},s(\widetilde{\nu}))=0$ holds, then, by the continuity of $\sigma(\widetilde{\nu},s)$ with respect to $s$, there is a ball :=${\mathcal{B}}:=\{s\in{\mathbb{R}}^{2}\;|\;\|s-s(\widetilde{\nu})\|\leq r\}$ with a positive radius $r$ centered at $s(\widetilde{\nu})$ such that $\sigma(\widetilde{\nu},s)\leq\epsilon$ for all $s\in{\mathcal{B}}$, which contradicts the optimality of $s(\widetilde{\nu})$ for (13) with $\nu=\widetilde{\nu}$. Consequently, $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$ is simple and nonzero. By continuity, there is an open neighborhood $\widetilde{U}\subseteq{\mathbb{R}}^{\mathsf{d}}$ of $\widetilde{\nu}$ and an open neighborhood $\widehat{U}\subseteq{\mathbb{R}}^{2}$ of $s(\widetilde{\nu})$ such that the singular value function $\sigma(\nu,s)$ remains simple and positive in $\widetilde{U}\times\widehat{U}$, while $p_{w}(\|s\|)>0$ in $\widehat{U}$, so $\sigma(\nu,s)$ as well as $g(s;\nu)$ defined as in (15) are real analytic in $\widetilde{U}\times\widehat{U}$ by Lemma 2.5. We assume, without loss of generality, that $g_{s}(s;\nu)\neq 0$, indeed the existence of a positive real number $\gamma$ such that $\|g_{s}(s;\nu)\|\geq\gamma$, for all $(s,\nu)\in\widetilde{U}\times\widehat{U}$ (if necessary by choosing $\widetilde{U}$, $\widehat{U}$ smaller).
The set of optimal solutions $S(\nu)$ of (13) is upper semicontinuous in $\widetilde{U}$ [3, Proposition 4.41] (i.e. for every neighborhood ${\mathcal{S}}$ of $S(\nu)$ there is a neighborhood ${\mathcal{V}}$ of $\nu\in\widetilde{U}$ such that ${\mathcal{S}}\supseteq S(\vartheta)$ for all $\vartheta\in{\mathcal{V}}$). In particular, there is a neighborhood $U\subseteq\widetilde{U}$ of $\widetilde{\nu}$ such that $S(\vartheta)\subseteq\widehat{U}$ for all $\vartheta\in U$. Let $s(\vartheta)$ denote a maximizer of (13) for $\vartheta\in U$. (It is soon to be established that (13) with $\vartheta\in U$ has a unique maximizer, so the notation $s(\vartheta)$ is consistent with the notation introduced before the theorem.) The singular value function $\sigma(\nu,s)$ is real analytic at $(\vartheta,s(\vartheta))\in U\times\widehat{U}\subseteq\widetilde{U}\times\widehat{U}$ and $g_{s}(s(\vartheta);\vartheta)\neq 0$ from the previous paragraph. The linear independence constraint qualification holds at $(\vartheta,s(\vartheta))$ due to $g_{s}(s(\vartheta);\vartheta)\neq 0$. Hence, by applying the first-order necessary conditions for a constrained optimization problem [30, Theorem 12.1] to (13) with $\nu=\vartheta$, there is a Lagrange multiplier $\mu(\vartheta)$ satisfying
[…]
It is evident from (18) that the Lagrange multiplier $\mu(\vartheta)$ corresponding to the maximizer $s(\vartheta)$ is unique. Letting, for $\nu\in U$, $y(\nu):=(s(\nu),\mu(\nu))$, indeed, the first-order necessary conditions imply ${\mathcal{L}}_{y}(\nu,y(\nu))=0$ for every $\nu\in U$. Note also that $s(\widetilde{\nu})$ must be the unique maximizer of (13) with $\nu=\widetilde{\nu}$.
On the other hand, since ${\mathcal{L}}_{yy}(\widetilde{\nu},y(\widetilde{\nu}))$ is invertible by the non-degeneracy of $\widetilde{\nu}$, the analytic implicit function theorem implies the existence of a neighborhood $U_{0}\subseteq U$ of $\widetilde{\nu}$ and a unique function $w:U_{0}\rightarrow\widehat{U}$ satisfying $w(\widetilde{\nu})=y(\widetilde{\nu})$ and ${\mathcal{L}}_{y}(\nu,w(\nu))=0$ for all $\nu\in U_{0}$. Moreover, the function $w$ is real analytic. By the uniqueness of $w$, the function $\nu\in U_{0}\mapsto y(\nu)$ is unique and $y(\nu)=w(\nu)$ for every $\nu\in U_{0}$. This also shows the real analyticity of $y(\nu)$ at every $\nu\in U_{0}$. In particular, the map $\nu\in U\mapsto\alpha_{\epsilon}(\nu)=s_{1}(\nu)$ is real analytic at $\nu=\widetilde{\nu}$.
To derive the formulas for the derivative given by (16) and (17), we note that
$\begin{split}&\sigma_{s_{1}}(\widetilde{\nu},s(\widetilde{\nu}))\;=\;\mathrm{Re}\left\{u^{\ast}\frac{\partial P}{\partial s_{1}}(\widetilde{\nu},s(\widetilde{\nu}))v\right\}\;=\;\mathrm{Re}\{u^{\ast}P^{\prime}(s(\widetilde{\nu});\widetilde{\nu})v\}\,,\\[1.99997pt] &\frac{\partial}{\partial s_{1}}\left[p_{w}(\|s\|)\right]\bigg{|}_{s=s(\widetilde{\nu})}\;=\;\frac{s_{1}(\widetilde{\nu})}{\|s(\widetilde{\nu})\|}\,p^{\prime}_{w}(\|s(\widetilde{\nu})\|)\;,\end{split}$
where, in the first line, we use the analytic formula in (14), in particular $u$, $v$ denote a pair of consistent unit left, right singular vectors corresponding to the singular value $\sigma(\widetilde{\nu},s(\widetilde{\nu}))$. It then follows from (18) that
$\mu(\widetilde{\nu})\;=\;\frac{1}{\mathrm{Re}\{u^{\ast}P^{\prime}(s(\widetilde{\nu});\widetilde{\nu})v\}-\epsilon\{s_{1}(\widetilde{\nu})/\|s(\widetilde{\nu})\|\}p^{\prime}_{w}(\|s(\widetilde{\nu})\|)}\,.$
Since $\mu(\nu)\neq 0$ for $\nu\in U_{0}$ due to (18), by the complementary conditions we must have $g(s(\nu);\nu)=0$ for $\nu\in U_{0}$ implying that $\alpha_{\epsilon}(\nu)=s_{1}(\nu)={\mathcal{L}}(\nu,y(\nu))$ for all $\nu\in U_{0}$. Finally, by also making use of ${\mathcal{L}}_{y}(\widetilde{\nu},y(\widetilde{\nu}))=0$, we deduce that
$\begin{split}\frac{\partial\alpha_{\epsilon}}{\partial\nu_{j}}(\widetilde{\nu})\;\;=\;\frac{\partial{\mathcal{L}}}{\partial\nu_{j}}(\widetilde{\nu},y(\widetilde{\nu}))+{\mathcal{L}}_{y}(\widetilde{\nu},y(\widetilde{\nu}))^{T}\frac{\partial y}{\partial\nu_{j}}(\widetilde{\nu})\;&=\;\frac{\partial{\mathcal{L}}}{\partial\nu_{j}}(\widetilde{\nu},y(\widetilde{\nu}))\\ &=\;-\mu(\widetilde{\nu})\frac{\partial\sigma}{\partial\nu_{j}}(\widetilde{\nu},s(\widetilde{\nu}))\\[3.00003pt] &=\;-\mu(\widetilde{\nu})\mathrm{Re}\left\{u^{\ast}P_{\nu_{j}}(\widetilde{\nu},s(\widetilde{\nu}))v\right\}\>,\end{split}$
where again the last equality follows from the formula in (14) for the derivatives of a singular value function.
Remark 2.8.
[…]
Remark 2.9.
Suppose that $P(\lambda;\nu)=\lambda^{2}M(\nu)+\lambda C(\nu)+K(\nu)$ is a Hermitian quadratic matrix polynomial for all $\nu\in\Omega$, i.e. $M(\nu),C(\nu),K(\nu)$ are Hermitian matrices for all $\nu\in\Omega$. Then it is apparent from the characterization in (5) that $\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ is symmetric in the complex plane with respect to the real axis, i.e. $z\in\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ if and only if $\,\overline{z}\in\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$, for all $\nu\in\Omega$.
As a result, the optimization problem in (13) for $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ can be restricted to $(s_{1},s_{2})\in{\mathbb{R}}^{2}$ with $s_{2}\geq 0$ rather than any $(s_{1},s_{2})\in{\mathbb{R}}^{2}$. The definition of a non-degenerate point $\nu\in\Omega$ can then be altered accordingly. In particular, now at a non-degenerate point $\nu$ the optimization problem is required to have a unique optimizer $s(\nu)=(s_{1}(\nu),s_{2}(\nu))$ in the upper half-plane $\{(s_{1},s_{2})\in{\mathbb{R}}^{2}\;|\;s_{2}\geq 0\}$ rather than ${\mathbb{R}}^{2}$. Theorem 2.6 in the Hermitian setting still holds with these changes.
3 Computation of the $\epsilon$-pseudospectral abscissa for large-scale problems
For the large matrix setting, inspired by the ideas in [14], a method for the computation of the $\epsilon$-pseudospectral abscissa of a nonlinear eigenvalue problem has been proposed in [20]. The approach finds a locally rightmost point in the $\epsilon$-pseudospectrum for a projected nonlinear eigenvalue problem by means of the support-function based optimization technique in [23], and refines the projected nonlinear eigenvalue problem by expanding the subspace based on the computed rightmost point.
Even through the approach in [20] can be adapted to our polynomial eigenvalue problem setting, it does not exploit the polynomial structure, in particular the tools such as the vertical and horizontal searches in the previous section. Additionally, the approach in [20] is developed for the slightly different definition of the $\epsilon$-pseudospectrum in [26, 32, 33] that uses the characterization $\{\lambda\in{\mathbb{C}}\;|\;\sigma_{\min}(P(\lambda))\leq\epsilon q_{w}(|\lambda|)\}$ for $q_{w}(z)=w_{m}z^{2}+w_{c}z+w_{k}$.
For these reasons we introduce a subspace framework for our setting, motivated by the approaches in [14, 20]. Given a subspace ${\mathcal{V}}$ of ${\mathbb{C}}^{\mathsf{n}}$ and a matrix $V$ whose columns form an orthonormal basis for this subspace, i.e. $\mathcal{V}=\mbox{\rm Col}(V)$, the subspace framework to compute $\alpha_{\epsilon}(P)$, replaces $P(\lambda)$ by the rectangular matrix polynomial
$P^{V}(\lambda)\;:=\;P(\lambda)V\;=\;\lambda^{2}(MV)+\lambda(CV)+(KV)\>.$
We define the associated $\epsilon$-pseudospectrum and the $\epsilon$-pseudospectral abscissa of the projected problem by
(19) $\begin{split}\Lambda^{\mathcal{V}}_{\epsilon}(P)&:=\;\{\lambda\in{\mathbb{C}}\;|\;\sigma_{\min}(P^{V}(\lambda))\leq\epsilon p_{w}(|\lambda|)\}\;,\;\;\text{and}\\[3.00003pt] \alpha^{\mathcal{V}}_{\epsilon}(P)&:=\;\sup\{\mathrm{Re}(z)\;|\;z\in{\mathbb{C}}\text{ s.t. }\sigma_{\min}(P^{V}(z))\leq\epsilon p_{w}(|z|)\}\;,\end{split}$
respectively. Clearly, the smallest singular value of $P^{V}(\lambda)$ depends on the subspace ${\mathcal{V}}$, yet is independent of the orthonormal basis $V$ that is used for ${\mathcal{V}}$. Hence, as it is made explicit by the notations, the definitions of $\Lambda^{\mathcal{V}}_{\epsilon}(P)$ and $\alpha^{\mathcal{V}}_{\epsilon}(P)$ are independent of the choice of the orthonormal basis $V$ for ${\mathcal{V}}$.
The matrix polynomial $P^{V}(\lambda)$ is of size $\mathsf{n}\times r$, where $r$ is the dimension of the subspace ${\mathcal{V}}$. The associated $\epsilon$-pseudospectrum and the $\epsilon$-pseudospectral abscissa can be equivalently defined in terms of a reduced matrix polynomial. In particular, let
$\left[\begin{array}[]{ccc}MV&CV&KV\end{array}\right]\;=\;Q\left[\begin{array}[]{ccc}\widehat{M}&\widehat{C}&\widehat{K}\end{array}\right]$
be a reduced QR factorization (see e.g. [9]) with $Q\in{\mathbb{C}}^{\mathsf{n}\times 3r}$ and $\widehat{M},\widehat{C},\widehat{K}\in{\mathbb{C}}^{3r\times r}$. Then for the matrix polynomial $\widehat{P}^{V}(\lambda)=\lambda^{2}\widehat{M}+\lambda\widehat{C}+\widehat{K}$ of size $3r\times r$, we have
$\sigma_{\min}(P^{V}(\lambda))\;=\;\sigma_{\min}(\widehat{P}^{V}(\lambda))\quad\mbox{\rm for all }\ \lambda\in{\mathbb{C}}.$
Hence, in the definitions of $\Lambda^{\mathcal{V}}_{\epsilon}(P)$ and $\alpha^{\mathcal{V}}_{\epsilon}(P)$ in (19) the $3\mathsf{n}\times r$ matrix polynomial $P^{V}$ can be replaced by the $3r\times r$ matrix polynomial $\widehat{P}^{V}$.
Algorithm 1 can be adapted to efficiently compute $\alpha^{\mathcal{V}}_{\epsilon}(P)$, as well as a rightmost point in $\Lambda^{\mathcal{V}}_{\epsilon}(P)$. The only modification is that in the vertical and horizontal searches, specifically in the definitions of ${\mathcal{M}}_{j}(x)$ and ${\mathcal{N}}_{j}(y)$ in Theorem 2.1 and Theorem 2.2, the matrix polynomial $P$ and matrix $M$ must be replaced by $\widehat{P}^{V}$ and $\widehat{M}$, respectively, and the sizes of the other zero, scalar-times-identity blocks of ${\mathcal{M}}_{j}(x)$, ${\mathcal{N}}_{j}(y)$ must be adjusted accordingly. This results in matrix polynomials ${\mathcal{M}}(\lambda;x)$ and ${\mathcal{N}}(\lambda;y)$ of size $4r\times 4r$ required by the reduced problems rather than their counterparts of size $2\mathsf{n}\times 2\mathsf{n}$ for the full problem.
The following properties of the $\epsilon$-pseudospectrum and $\epsilon$-pseudospectral abscissa have been shown in Lemma 4.1 and Theorem 4.3 of [20].
Theorem 3.1 (Monotonicity).
Let $\,{\mathcal{U}}$, ${\mathcal{V}}$ be two subspaces of $\,{\mathbb{C}}^{n}$ such that $\,{\mathcal{U}}\subseteq{\mathcal{V}}$. The following assertions hold:
  1. 1.
    $\Lambda^{\mathcal{U}}_{\epsilon}(P)\;\subseteq\;\Lambda^{\mathcal{V}}_{\epsilon}(P)\;\subseteq\;\Lambda_{\epsilon}(P)$;
  2. 2.
    $\alpha^{\mathcal{U}}_{\epsilon}(P)\;\leq\;\alpha^{\mathcal{V}}_{\epsilon}(P)\;\leq\;\alpha_{\epsilon}(P)$.
Theorem 3.2 (Low Dimensionality).
The following assertions are equivalent for a subspace ${\mathcal{V}}$ of ${\mathbb{C}}^{n}$:
  1. 1.
    $\alpha_{\epsilon}(P)\;=\;\alpha^{\mathcal{V}}_{\epsilon}(P)$;
  2. 2.
    The subspace ${\mathcal{V}}$ contains a vector $v$ that is a right singular vector corresponding to $\sigma_{\min}(P(\underline{z}))$ for some $\underline{z}\in\Lambda_{\epsilon}(P)$ such that $\mathrm{Re}(\underline{z})=\alpha_{\epsilon}(P)$.
The subspace framework to compute $\alpha_{\epsilon}(P)$ and a rightmost point $\underline{z}$ in $\Lambda_{\epsilon}(P)$ when $\mathsf{n}$ is large is presented in Algorithm 2 below. Note that throughout this description ${\mathcal{V}}_{k}$ is the restriction subspace spanned by the columns of the matrix $V_{k}$. The subspace framework gradually expands the subspace ${\mathcal{V}}_{k}$ until the correct vector or a nearby vector as in Theorem 3.2 is in this subspace, ensuring that $\alpha_{\epsilon}(P)\approx\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(P)$. At every iteration, a rightmost point $z^{(k)}$ in the reduced pseudospectrum is computed by means of Algorithm 1 in line 4. If the condition for termination is not met, then the subspace is expanded by including a right singular vector of $P(z^{(k)})$ corresponding to its smallest singular value in line 6.
0:  Quadratic matrix polynomial $P(\lambda)$ of the form (3) with large $\mathsf{n}$, and $\epsilon>0$.
0:  Estimates $\underline{x}$ for $\alpha_{\epsilon}(P)$, and $\underline{z}\in{\mathbb{C}}$ for a rightmost point in $\Lambda_{\epsilon}(P)$.
1:  $z^{(0)}\leftarrow$ an eigenvalue of $P(\lambda)$.
2:  $V_{0}\leftarrow$ a unit norm right singular vector corresponding to $\sigma_{\min}(P(z^{(0)}))$.
3:  for $k=1,2,\dots$ do
4:     Solve Reduced Problem. Find a rightmost point $z^{(k)}$ in $\Lambda^{{\mathcal{V}}_{k-1}}_{\epsilon}(P)$ using Algorithm 1.
5:     Return $\underline{x}\leftarrow\text{Re}(z^{(k)})\,$ and $\,\underline{z}\leftarrow z^{(k)}\;$ if convergence occurred.
6:     Expand Subspace. Compute a right singular vector $v$ corresponding to $\sigma_{\min}(P(z^{(k)}))$, and set $V_{k}\leftarrow\text{orth}\left(\left[\begin{array}[]{cc}V_{k-1}&v\end{array}\right]\right)$.
7:  end for
Algorithm 2 Subspace framework for large-scale computation of $\alpha_{\epsilon}(P)$
Convergence. A variant of the subspace framework for the matrix setting in [14] converges to a locally rightmost point in the $\epsilon$-pseudospectrum. We observe a similar local convergence behavior for Algorithm 2 in the quadratic matrix polynomial setting. The choice of the initial projection subspace has an important role in attaining global convergence. To this end, the initialization in line 2 can possibly be done in a more sophisticated way. Additionally, it has been shown in [14] that $\text{Re}(z^{(k)})$ converges as $k\rightarrow\infty$ rather quickly at a superlinear rate. Such a quick convergence is expected in our polynomial eigenvalue setting as well due to the interpolation properties that are indicated next.
Interpolation. Suppose that the computed rightmost point $z^{(\ell)}$ after the execution of line 4 at the $\ell$th iteration, $\ell\geq 1$, is such that $\sigma_{\min}(P(z^{(\ell)}))$ is simple and nonzero. Using the variational characterization of the smallest singular value [9], as well as the analytical formulas for the derivatives of singular value functions (in particular Lemma 2.5), it can be shown for Algorithm 2 that the Hermite interpolation properties
(20) $\begin{split}&\sigma_{\min}(P(z^{(\ell)}))=\sigma_{\min}(P^{V_{k}}(z^{(\ell)}))\,,\\[1.99997pt] &\frac{\partial}{\partial\,\mathrm{Re}(z)}[\sigma_{\min}(P(z))]\bigg{|}_{z=z^{(\ell)}}=\frac{\partial}{\partial\,\mathrm{Re}(z)}[\sigma_{\min}(P^{V_{k}}(z))]\bigg{|}_{z=z^{(\ell)}}\,,\\[1.99997pt] &\frac{\partial}{\partial\,\mathrm{Im}(z)}[\sigma_{\min}(P(z))]\bigg{|}_{z=z^{(\ell)}}=\frac{\partial}{\partial\,\mathrm{Im}(z)}[\sigma_{\min}(P^{V_{k}}(z))]\bigg{|}_{z=z^{(\ell)}}\,\end{split}$
are satisfied for every $k\geq\ell$. Note that the interpolation of the singular values $\sigma_{\min}(P(z^{(\ell)}))=\sigma_{\min}(P^{V_{k}}(z^{(\ell)}))$ holds even if $\sigma_{\min}(P(z^{(\ell)}))$ is not simple or is zero.
Termination. In practice, we terminate the algorithm in line 5 when the change in $\underline{x}$ in two consecutive iterations is less than a tolerance in a relative sense.
Large-scale requirements and computational cost. The only large-scale problems whose solutions are required by Algorithm 2 are the computation of an eigenvalue of $P(\lambda)$ for initialization, and the singular vector computation corresponding to $\sigma_{\min}(P(z^{(0)}))$ in line 2, as well as corresponding to $\sigma_{\min}(P(z^{(k)}))$ in line 6 at every iteration. Both of these computations can be carried out using iterative methods such as an appropriate Krylov subspace method [16], e.g. Arnoldi or Lanczos method. In practice, we use eigs, an implementation of an implicitly restarted Arnoldi method, in MATLAB. Krylov subspace methods are especially well-suited for large sparse eigenvalue and singular value problems, even though they may converge to eigenvalues, singular values other than the targeted ones, or exhibit very slow convergence in some cases. Clearly, reliability of Algorithm 2 relies on the reliability of Krylov subspace methods. In our setting here, they usually work quite effectively and reliably especially on problems arising from damping optimization. The most expensive computational task at every iteration of Algorithm 2 is the smallest singular value computation in line 6. By using a Krylov subspace method, this usually has computational cost ${\mathcal{O}}(\mathsf{n}^{2})$ (possibly ${\mathcal{O}}(\mathsf{n})$ if the matrix polynomial is sparse), but it may be higher especially if the convergence of the Krylov subspace method is slow. Assuming a superlinear convergence of Algorithm 2, the Krylov subspace method to compute the smallest singular value in line 6 is called only a few times. As a result, it is reasonable to assume that the overall cost of Algorithm 2 is ${\mathcal{O}}(\mathsf{n}^{2})$, but possibly with a big constant hidden in the ${\mathcal{O}}$-notation.
4 Large-Scale minimization of the $\epsilon$-pseudospectral abscissa
In this section, we discuss the minimization of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ over $\nu\in\underline{\Omega}$ in the large-scale setting, i.e. when $\mathsf{n}$ is large. Similar to the ideas in the previous section, but now for minimization and not for computation, the large parameter-dependent matrix polynomial $P(\,\lambda\,;\nu)$ is replaced by
$P^{V}(\lambda;\nu)\;:=\;P(\lambda;\nu)V\;=\;\lambda^{2}M^{V}(\nu)+\lambda C^{V}(\nu)+K^{V}(\nu)\,,$
where
$\begin{split}&M^{V}(\nu)\;\;:=\;\;M(\nu)V\;=\;\>\psi_{1}(\nu)(M_{1}V)+\dots+\psi_{\mathsf{m}}(\nu)(M_{\mathsf{m}}V)\,,\\[3.00003pt] &\,C^{V}(\nu)\;\;:=\;\;C(\nu)V\;\>=\;\>\zeta_{1}(\nu)(C_{1}V)+\dots+\zeta_{\mathsf{c}}(\nu)(C_{\mathsf{c}}V)\,,\\[3.00003pt] &\,K^{V}(\nu)\;\;:=\;\;K(\nu)V\;\,=\;\>\kappa_{1}(\nu)(K_{1}V)+\dots+\kappa_{\mathsf{k}}(\nu)(K_{\mathsf{k}}V)\,\end{split}$
for a matrix $V\in{\mathbb{C}}^{\mathsf{n}\times r}$ whose columns form an orthonormal basis for a small $r\ll\mathsf{n}$ dimensional restriction subspace ${\mathcal{V}}$. We form $M_{i}V,\,C_{j}V,\,K_{\ell}V\in{\mathbb{C}}^{\mathsf{n}\times r}$ for $i=1,\dots,\mathsf{m}$, $j=1,\dots,\mathsf{c}$, $\>\ell=1,\dots,\mathsf{k}\,$ in advance, then minimize
(21) $\alpha^{\mathcal{V}}_{\epsilon}(P(\,\cdot\,;\nu))\;:=\;\sup\{\mathrm{Re}(z)\;|\;z\in{\mathbb{C}}\text{ s.t. }\sigma_{\min}(P^{V}(z;\nu))\leq\epsilon p_{w}(|z|)\}$
over $\nu\in\underline{\Omega}$, which we refer to as a reduced problem. The value of $\alpha^{\mathcal{V}}_{\epsilon}(P(\,\cdot\,,\nu))$ at a given $\nu$, as well as the corresponding $z$ maximizing the right-hand side in (21) can be computed efficiently using a variant of Algorithm 1 as discussed in Section 3.
We carry out the minimization of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ over $\nu\in\underline{\Omega}$ in the large-scale setting using the subspace restriction idea as follows. At every iteration, we solve a reduced $\epsilon$-pseudospectral abscissa minimization problem. Then a right-most point in $\Lambda_{\epsilon}(P(:,\nu))$ is computed at the minimizing parameter value for the reduced problem. Finally, we enrich the restriction subspace by a right singular vector corresponding to $\sigma_{\min}(P(z;\nu))$ at the minimizing parameter value $\nu$ and the computed right-most point $z$.
This subspace approach is formally presented in Algorithm 3, with ${\mathcal{V}}_{k}:=\text{Col}(V_{k})$, and the iterates $\nu^{(k)},z^{(k)}$ at the $k$th step.
0:  Quadratic matrix polynomial $P(\lambda;\nu)$ of the form (1) with large $\mathsf{n}$, the feasible region $\underline{\Omega}\,$, and $\epsilon>0$.
0:  Estimates $\underline{\nu}$ for $\arg\min_{\nu\in\underline{\Omega}}\alpha_{\epsilon}(P(\,\cdot\,;\nu))$, and $\underline{z}\in{\mathbb{C}}$ for a globally rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\underline{\nu}))$.
1:  Initialize. $\nu^{(0)}_{1},\dots\nu^{(0)}_{\eta}\leftarrow$ initially chosen points in $\underline{\Omega}$. $z^{(0)}_{j}\leftarrow$ a rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\nu^{(0)}_{j}))$ using Algorithm 2 for $j=1,\dots,\eta$. $v^{(0)}_{j}\leftarrow$ a right singular vector corresponding to $\sigma_{\min}(P(z^{(0)}_{j};\nu^{(0)}_{j}))$ for $j=1,\dots,\eta$. $V_{0}\leftarrow$ a matrix whose columns form an orthonormal basis for span$\{v^{(0)}_{1},\dots v^{(0)}_{\eta}\}$.
2:  for $k=1,2,\dots$ do
3:     Solve Reduced Minimization Problem. $\nu^{(k)}\leftarrow\arg\min_{\nu\in\underline{\Omega}}\alpha_{\epsilon}^{{\mathcal{V}}_{k-1}}(P(\,\cdot\,;\nu))$.
4:     $z^{(k)}\leftarrow$ a rightmost point in $\Lambda_{\epsilon}(P(\,\cdot\,;\nu^{(k)}))$ using Algorithm 2.
5:     Return $\underline{\nu}\leftarrow\nu^{(k)}$, $\underline{z}\leftarrow z^{(k)}$ if convergence occurred.
6:     Expand Subspace. Compute a right singular vector $v^{(k)}$ corresponding to $\sigma_{\min}(P(z^{(k)};\nu^{(k)}))$, and set $V_{k}\leftarrow\text{orth}\left(\left[\begin{array}[]{cc}V_{k-1}&v^{(k)}\end{array}\right]\right)$.
7:  end for
Algorithm 3 Subspace framework for large-scale minimization of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$
In practice, we terminate Algorithm 3 in line 5 when the optimal values of the reduced minimization problems in two consecutive iterations do not differ by more than a prescribed tolerance. A remarkable feature of Algorithm 3 is that it does not require the computation of all eigenvalues of large matrix polynomials. The only required large-scale computations are the computation of the right singular vector corresponding to the smallest singular value of large matrices in lines 1, 6, and in Algorithm 2, as well as the computation of an eigenvalue of a large matrix polynomial to initialize Algorithm 2. In practice, we perform these smallest singular value and eigenvalue computations by means of Krylov subspace methods as noted at the end of Section 3. The limitations of the Krylov subspace methods listed there apply, but in practice we usually observe good convergence behavior which is due to the fact that in the context of the optimization we usually have very good starting values.
The main ingredients that determine the runtime of Algorithm 3 are
  1. (i)
    the solution of the reduced minimization problem in line 3,
  2. (ii)
    the $\epsilon$-pseudospectral abscissa computation on the full problem via Algorithm 2 in line 4,
  3. (iii)
    the smallest singular value computation on the full problem in line 6
at every iteration. As argued in Section 3, it is reasonable to assume both (ii) and (iii) have ${\mathcal{O}}(\mathsf{n}^{2})$ computational cost. However, Algorithm 2 itself computes the smallest singular value a few times, so (ii) is usually more expensive than (iii). Thus, the overall cost of Algorithm 3 is mainly determined by the cost of (ii) and (iii) times the number of iterations of Algorithm 3. For large $\mathsf{n}$, we expect (ii) to dominate, as (i) involves only small reduced problems, that is the overall computational cost may largely be determined by the number of calls to Algorithm 2.
4.1 Solution of reduced minimization problems
Just as for the full $\epsilon$-pseudospectral abscissa function, the objective to be minimized $\alpha^{{\mathcal{V}}}_{\epsilon}(\nu):=\alpha^{{\mathcal{V}}}_{\epsilon}(P(\,\cdot\,;\nu))$ for a reduced minimization problem is nonconvex and nonsmooth. Thus, $\alpha^{{\mathcal{V}}}_{\epsilon}(\nu)$ can again be minimized globally using EigOpt if there are only a few parameters, or locally using GRANSO if there are many parameters. We again use the variant of Algorithm 1 discussed in Section 3 to evaluate the objective function $\alpha^{{\mathcal{V}}}_{\epsilon}(\nu)$, and to determine a right-most point in $\Lambda^{{\mathcal{V}}}_{\epsilon}(P(\,\cdot\,;\nu))$ from which the gradient of $\alpha^{{\mathcal{V}}}_{\epsilon}(\nu)$ can be obtained as shown next. The arguments here are analogous to those for the full $\epsilon$-pseudospectral abscissa function.
The reduced $\epsilon$-pseudospectral abscissa function $\alpha^{{\mathcal{V}}}_{\epsilon}(\nu)$ can be written as
(22) $\alpha^{\mathcal{V}}_{\epsilon}(\nu)\;=\;\max\{s_{1}\;|\;s=(s_{1},s_{2})\in{\mathbb{R}}^{2}\;\text{ is s.t. }\;\sigma^{\mathcal{V}}(\nu,s)-\epsilon p_{w}(\|s\|)\leq 0\}\,,$
where $\sigma^{\mathcal{V}}(\nu,s):=\sigma_{\min}(P^{V}(\nu,s))$, $P^{V}(\nu,s):=P^{V}(s_{1}+{\rm i}s_{2};\nu)$ for $s=(s_{1},s_{2})\in{\mathbb{R}}^{2}$ and for any matrix $V$ whose columns form an orthonormal basis for ${\mathcal{V}}$. The associated Lagrangian function is
${\mathcal{L}}^{\mathcal{V}}(\nu,s,\mu)\;:=\;s_{1}-\mu\{\sigma^{\mathcal{V}}(\nu,s)-\epsilon p_{w}(\|s\|)\}\;.$
which we express as ${\mathcal{L}}^{\mathcal{V}}(\nu,y)$ where again $y=(s,\mu)$. We reserve the notations ${\mathcal{L}}^{\mathcal{V}}_{y}$ and ${\mathcal{L}}^{\mathcal{V}}_{yy}$ for the gradient and the Hessian of ${\mathcal{L}}^{\mathcal{V}}$ with respect to $y$. Assuming the uniqueness of the maximizer $s^{\mathcal{V}}(\nu)$ for the maximization problem in (22), the simplicity of the singular value $\sigma^{\mathcal{V}}(\nu,s^{\mathcal{V}}(\nu))$, and the satisfaction of a constraint qualification for (22) at $s=s^{\mathcal{V}}(\nu)$, there is a unique $\mu^{\mathcal{V}}(\nu)$ satisfying ${\mathcal{L}}^{\mathcal{V}}_{y}(\nu,s^{\mathcal{V}}(\nu),\mu^{\mathcal{V}}(\nu))=0$. Similar to the practice before, we use $y^{\mathcal{V}}(\nu):=(s^{\mathcal{V}}(\nu),\mu^{\mathcal{V}}(\nu))$.
As in the non-reduced case, a point $\nu\in\Omega$ is called a non-degenerate point of the restriction of $P(\,\cdot\,;\,\cdot\,)$ to a subspace ${\mathcal{V}}$ if there is a unique optimizer of (22), denoted as $s^{\mathcal{V}}(\nu)=(s^{\mathcal{V}}_{1}(\nu),s^{\mathcal{V}}_{2}(\nu))$, the singular value $\sigma^{\mathcal{V}}(\nu,s^{\mathcal{V}}(\nu))$ is simple, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{\mathcal{V}}(\nu,y^{\mathcal{V}}(\nu))$ is invertible.
The reduced problem satisfies an analogue of Theorem 2.6 but now by operating on ${\mathcal{L}}^{\mathcal{V}}(\nu,y)$ rather than ${\mathcal{L}}(\nu,y)$. Similar to the requirement in Theorem 2.6 on the gradient $g_{s}$ of $g$ defined as in (15), we need a requirement here but now in terms of
$g^{\mathcal{V}}(s;\nu)\;:=\;\sigma^{\mathcal{V}}(\nu,s)-\epsilon p_{w}(\|s\|),$
in particular in terms of its gradient with respect to $s$, which we denote by $g^{\mathcal{V}}_{s}$.
Theorem 4.1.
Let ${\mathcal{V}}$ be a subspace of ${\mathbb{C}}^{n}$, and $V$ be any matrix whose columns form an orthonormal basis for ${\mathcal{V}}$. Moreover, let $\,\widetilde{\nu}$ be a non-degenerate point for the restriction of $P(\,\cdot\,;\,\cdot\,)$ to ${\mathcal{V}}$, and suppose that $p_{w}(\|s^{\mathcal{V}}(\widetilde{\nu})\|)\neq 0\,$, $\,g^{\mathcal{V}}_{s}(s^{\mathcal{V}}(\widetilde{\nu});\widetilde{\nu})\neq 0$. The function $\nu\mapsto\alpha^{\mathcal{V}}_{\epsilon}(\nu)$ is real analytic at $\widetilde{\nu}$ with the first derivatives
(23) $\frac{\partial\alpha^{\mathcal{V}}_{\epsilon}}{\partial\nu_{j}}(\widetilde{\nu})\;=\;-\mu^{\mathcal{V}}(\widetilde{\nu})\mathrm{Re}\left\{u^{\ast}P^{V}_{\nu_{j}}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}))v\right\}\,,\;\;\;\;j=1,\dots,\mathsf{d}\,,$
where $P^{V}_{\nu_{j}}$ denotes the partial derivative of $P^{V}$ with respect to $\nu_{j}$,
(24) $\mu^{\mathcal{V}}(\widetilde{\nu})\;:=\;\frac{1}{\mathrm{Re}\{u^{\ast}P^{\prime}(s^{\mathcal{V}}(\widetilde{\nu});\widetilde{\nu})Vv\}-\epsilon\{s^{\mathcal{V}}_{1}(\widetilde{\nu})/\|s^{\mathcal{V}}(\widetilde{\nu})\|\}p^{\prime}_{w}(\|s^{\mathcal{V}}(\widetilde{\nu})\|)}\,,$
and $u,v$ are consistent unit left, right singular vectors, respectively, corresponding to the smallest singular value $\sigma^{\mathcal{V}}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}))$ of $P^{V}(\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu}))$.
Proof 4.2.
The proof is similar to that for Theorem 2.6.
4.2 Interpolation properties
We now show that the subspace framework in Algorithm 3 is interpolatory to the full problem. Using the abbreviations $\Lambda_{\epsilon}(\nu):=\Lambda_{\epsilon}(P(\,\cdot\,;\nu))$ and $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu):=\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(P(\,\cdot\,;\nu))$, we have the following result.
Theorem 4.3 (Hermite Interpolation).
Consider the subspaces and sequences generated by Algorithm 3. Then the following assertions hold for every $k\geq 1$ and $\ell=1,\dots,k$:
  1. 1.
    $\alpha_{\epsilon}(\nu^{(\ell)})=\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$;
  2. 2.
    The point $z^{(\ell)}$ is a rightmost point in $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$;
  3. 3.
    Suppose $\nu^{(\ell)}$ is a non-degenerate point such that $p_{w}(\|\widehat{z}^{(\ell)}\|)\neq 0$, $g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$ for $\widehat{z}^{(\ell)}:=(\text{Re}(z^{(\ell)}),\text{Im}(z^{(\ell)}))$, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{{\mathcal{V}}_{k}}(\nu^{(\ell)},\widehat{z}^{(\ell)},\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)}))$ is invertible, where $\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)})$ is as in (24) but with $\nu^{(\ell)},\widehat{z}^{(\ell)}$ replacing $\widetilde{\nu},s^{\mathcal{V}}(\widetilde{\nu})$. Then $\alpha_{\epsilon}(\nu)$, $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ are differentiable at $\nu=\nu^{(\ell)}$, and $\nabla\alpha_{\epsilon}(\nu^{(\ell)})=\nabla\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$.
Proof 4.4.
1. The subspace ${\mathcal{V}}_{k}$ contains a right singular vector $v^{(\ell)}$ corresponding to $\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))$, where $z^{(\ell)}$ is a rightmost point in $\Lambda_{\epsilon}(\nu^{(\ell)})$. Hence, the assertion $\alpha_{\epsilon}(\nu^{(\ell)})=\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$ follows from Theorem 3.2.
2. By Theorem 3.1, part 2., we have
(25) $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})\;\leq\;\alpha_{\epsilon}(\nu^{(\ell)})=\mathrm{Re}(z^{(\ell)})\,.$
Moreover, as $v^{(\ell)}\in{\mathcal{V}}_{k}$, there exists a vector $a$ such that $v^{(\ell)}=V_{k}a$. Combining this with $z^{(\ell)}\in\Lambda_{\epsilon}(\nu^{(\ell)})$, we obtain
$\begin{split}\epsilon\;\geq\;\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;&=\;\frac{\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|}{p_{w}(|z^{(\ell)}|)}\\ &\geq\;\frac{\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\,,\end{split}$
[…]
(26) $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})\;\geq\;\alpha_{\epsilon}(\nu^{(\ell)})=\mathrm{Re}(z^{(\ell)})\,.$
Combining (25) and (26), we deduce that $z^{(\ell)}\in\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$ satisfies $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})=\mathrm{Re}(z^{(\ell)})$ as claimed.
3. As $\nu^{(\ell)}$ is a non-degenerate point, and $g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$ for $\widehat{z}^{(\ell)}:=(\text{Re}(z^{(\ell)}),\text{Im}(z^{(\ell)}))$, it follows from Theorem 2.6 that $\alpha_{\epsilon}(\nu)$ is differentiable at $\nu=\nu^{(\ell)}$.
We claim that $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ is also differentiable at $\nu=\nu^{(\ell)}$. To show this, recall by part 2. that $z^{(\ell)}$ is a rightmost point of $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$. Suppose there is another rightmost point $\widetilde{z}$ of $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$ satisfying $\widetilde{z}\in\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$ and
$\mathrm{Re}(\widetilde{z})=\mathrm{Re}(z^{(\ell)})=\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})=\alpha_{\epsilon}(\nu^{(\ell)}).$
Due to Theorem 3.1, part 1., we have $\widetilde{z}\in\Lambda_{\epsilon}(\nu^{(\ell)})$. Hence, $\widetilde{z}$ is also a rightmost point of $\Lambda_{\epsilon}(\nu^{(\ell)})$ in addition to $z^{(\ell)}$, a contradiction. This shows that $z^{(\ell)}$ is the unique rightmost point of $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$ as well.
For the proof of differentiability of $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ at $\nu=\nu^{(\ell)}$, we also need to show that $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ is simple. For the sake of contradiction, suppose otherwise that this singular value is not simple. Then it follows from the first-order optimality conditions, in particular from the analogue of (18) where $\nu^{(\ell)}$ replaces $\vartheta$, that the Lagrange multiplier $\mu(\nu^{(\ell)})$ is nonzero. Thus, by the complementarity condition $\mu(\nu^{\ell})g(z^{(\ell)};\nu^{(\ell)})=0$ [30, Theorem 12.1], as $z^{(\ell)}$ is the maximizer of the optimization problem (13) with $\nu=\nu^{(\ell)}$, we have $g(z^{(\ell)};\nu^{(\ell)})=0$ or equivalently $\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}=\epsilon$. Combining this with
$\begin{split}\epsilon\;\geq\;\frac{\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;&=\;\frac{\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|}{p_{w}(|z^{(\ell)}|)}\;\geq\;\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\,,\end{split}$
where $a$ is a unit norm right singular vector corresponding to $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$, we deduce that
$\frac{\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;=\;\frac{\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))}{p_{w}(|z^{(\ell)}|)}\;=\;\epsilon\,.$
Since $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ is not simple, there is another unit right singular vector $\widetilde{a}$ corresponding to $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ such that $a$ and $\widetilde{a}$ are orthogonal. But then
$\begin{split}\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))\;&=\;\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))\\ \;&=\;\|P(z^{(\ell)};\nu^{(\ell)})V_{k}a\|\;=\;\|P(z^{(\ell)};\nu^{(\ell)})V_{k}\widetilde{a}\|\,,\end{split}$
yielding the contradiction that $V_{k}a$, $V_{k}\widetilde{a}$ are orthogonal, right singular vectors corresponding to $\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))$, that is $\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))$ is not simple.
As $z^{(\ell)}$ is the unique rightmost point in $\Lambda^{{\mathcal{V}}_{k}}_{\epsilon}(\nu^{(\ell)})$, the smallest singular value $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$ is simple, and the Hessian $\nabla^{2}_{yy}{\mathcal{L}}^{{\mathcal{V}}_{k}}(\nu^{(\ell)},\widehat{z}^{(\ell)},\mu^{{\mathcal{V}}_{k}}(\nu^{(\ell)}))$ is invertible by assumption, $\nu^{(\ell)}$ is a non-degenerate point for the restriction of $P(\cdot;\cdot)$ to ${\mathcal{V}}_{k}$. Additionally, by the interpolation of the derivatives in (20), we have $g^{{\mathcal{V}}_{k}}_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})=g_{s}(\widehat{z}^{(\ell)};\nu^{(\ell)})\neq 0$. Consequently, $\alpha^{{\mathcal{V}}_{k}}_{\epsilon}(\nu)$ is differentiable at $\nu=\nu^{(\ell)}$ by Theorem 4.1.
Finally, for the equality of the gradients, suppose that $u$, $v$ are consistent unit norm left, right singular vectors, respectively, corresponding to $\sigma:=\sigma_{\min}(P(z^{(\ell)};\nu^{(\ell)}))$. Due to the simplicity assumption on this singular value, we have $v\in{\mathcal{V}}_{k}$. As a consequence, there exists a unit norm vector $a$ such that $v=V_{k}a$. Since $u$, $v$ are consistent left, right singular vectors, respectively, we have $\,P(z^{(\ell)};\nu^{(\ell)})v\,=\,\sigma u\,$ and $\,u^{\ast}P(z^{(\ell)};\nu^{(\ell)})\,=\,\sigma v^{\ast}\,$, which can be rewritten in terms of $a$ as
$P^{V_{k}}(z^{(\ell)};\nu^{(\ell)})a\;=\;\sigma u\quad\text{and}\quad u^{\ast}P^{V_{k}}(z^{(\ell)};\nu^{(\ell)})\;=\;\sigma a^{\ast}\,.$
Hence, $u$, $a$ are consistent unit left, right singular vectors, respectively, corresponding to $\sigma_{\min}(P^{V_{k}}(z^{(\ell)};\nu^{(\ell)}))$. From the expressions (16), (17) for the derivatives of the pseudospectral abscissa function, and (23), (24) for the reduced counterparts, we obtain
$\begin{split}\frac{\partial\alpha_{\epsilon}}{\partial\nu_{j}}(\nu^{(\ell)})\;&=\;\frac{-\mathrm{Re}\left\{u^{\ast}P_{\nu_{j}}(z^{(\ell)};\nu^{(\ell)})v\right\}}{\mathrm{Re}\{u^{\ast}P^{\prime}(z^{(\ell)};\nu^{(\ell)})v\}-\epsilon\{\mathrm{Re}(z^{(\ell)})/|z^{(\ell)}|\}p^{\prime}_{w}(|z^{(\ell)}|)}\\[3.00003pt] &=\;\frac{-\mathrm{Re}\left\{u^{\ast}P^{V_{k}}_{\nu_{j}}(z^{(\ell)};\nu^{(\ell)})a\right\}}{\mathrm{Re}\{u^{\ast}P^{\prime}(z^{(\ell)};\nu^{(\ell)})V_{k}a\}-\epsilon\{\mathrm{Re}(z^{(\ell)})/|z^{(\ell)}|\}p^{\prime}_{w}(|z^{(\ell)}|)}\\[6.99997pt] &=\;\frac{\partial\alpha^{{\mathcal{V}}_{k}}_{\epsilon}}{\partial\nu_{j}}(\nu^{(\ell)})\end{split}$
for $j=1,\dots,\mathsf{d}$, where $P_{\nu_{j}}(z;\nu)$, $P^{V_{k}}_{\nu_{j}}(z;\nu)$ denote the partial derivatives of $P$, $P^{V_{k}}$, respectively, with respect $\nu_{j}$ keeping $z$ and the entries of $\nu$ other than $\nu_{j}$ fixed.
4.3 Global convergence
The result below concerns the convergence of Algorithm 3, and is identical to [2, Theorem 3.1]. This result follows from part 2. of Theorem 3.1 and part 1. of Theorem 4.3.
Theorem 4.5.
Suppose that the iterates $\nu^{(\ell)}$ and $\nu^{(k)}$ by Algorithm 3 are equal for some integers $\ell$, $k$ such that $1\leq\ell<k$. Then $\nu^{(\ell)}$ is a global minimizer of $\alpha_{\epsilon}(\nu)$ over all $\nu\in\underline{\Omega}$.
Proof 4.6.
The proof is the same as that of [2, Theorem 3.1].
As we deal with finite dimensional $\mathsf{n}\times\mathsf{n}$ matrix polynomials, the restriction subspace ${\mathcal{V}}_{k}$ in Algorithm 3 must stagnate after finitely many iterations, and Theorem 4.5 implies that when stagnation in the subspaces occurs, then the iterates stagnate at a global minimizer.
After deriving the theoretical basis of the subspace framework, in the next section we apply the results to damping optimization problems.
5 Application to damping optimization
The classical damping optimization problem (see, e.g. [34]) concerns a second-order system of the form
(27) $M\ddot{x}(t)+C(\nu)\dot{x}(t)+Kx(t)\;=\;f(t)\,,$
where $M,K\in{\mathbb{R}}^{n\times n}$ are the symmetric positive definite mass, stiffness matrices, and $C(\nu)\in{\mathbb{R}}^{n\times n}$ is the symmetric positive semi-definite damping matrix dependent on parameters, e.g. the viscosities of passive dampers. Typically, $C(\nu)$ can be partitioned as
$C(\nu)\;=\;C_{\text{int}}+C_{\text{ext}}(\nu)$
with $C_{\text{int}}\;=\;2\xi M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2}$ representing the internal damping for some constant $\xi>0$, and $C_{\text{ext}}(\nu)$ representing the external damping that is determined by the viscosities $\nu$, as well as the positions of the dampers.
There are several objectives considered in the literature to choose the viscosities [27, 36]. One of the objectives that has been taken into consideration is the minimization of the spectral abscissa [7, 28]. In order to obtain robust stability, rather than minimizing the spectral abscissa, here we aim to choose the viscosities $\nu\in\underline{\Omega}$ minimizing the $\epsilon$-pseudospectral abscissa $\alpha_{\epsilon}(P(\cdot;\nu))$ for a given $\epsilon>0$, where $P(\lambda;\nu)=\lambda^{2}M+\lambda C(\nu)+K$, and $\underline{\Omega}$ is a prescribed box in the parameter space. In the last example we consider, in addition to a damper, that the matrix $K$ also depends on a parameter, and we still minimize $\alpha_{\epsilon}(P(\cdot;\nu))$ on a box $\underline{\Omega}$. In all of the examples in this section, by construction, the coefficients $M$, $C(\nu)$, $K$ are symmetric, and positive definite, positive semi-definite, positive definite, respectively.
All of the numerical experiments in this section were performed in MATLAB 2024b on a MacBook Air with Mac OS 15.4 operating system, Apple M3 CPU and 16GB RAM using our implementations of Algorithm 3 and 2, as well as other ingredients such as the criss-cross algorithm (i.e., Algorithm 1). The reduced minimization problem in line 3 of Algorithm 3 is solved globally throughout by means of the package EigOpt. When using Algorithm 3, the number of initial interpolation points in line 1 is $\eta=5$ and $\eta=8$ when the problem is dependent on one and two parameters, respectively. The initial interpolation points are chosen as equally spaced points on the interval $\underline{\Omega}$ if there is one parameter or on the line segment that connects the lower-left and upper-right corners of the box $\underline{\Omega}$ if there are two parameters. The prescribed termination tolerance for Algorithm 3 for the examples below is $10^{-8}$.
The MATLAB software with which we have carried out the numerical experiments in this section is publicly available [21]. All of the results reported below can be reproduced using this software.
Example 5.1.
This is a small-scale example arising from a mass-spring-damper system consisting of four masses, where the damper is positioned on the second mass, and we want to determine the optimal viscosity for this damper minimizing the $\epsilon$-pseudospectral abscissa objective. The system has the coefficients
$M=\mathrm{diag}(1,2,3,4)\>,\;\;C(\nu)=C_{\mathrm{int}}\,+\,\nu e_{2}e_{2}^{T}\>,\;\;K=\mathrm{tridiag}(-5,10,-5)\,,$
where $C_{\mathrm{int}}\,=\,\xi 2M^{1/2}\sqrt{M^{-1/2}KM^{-1/2}}M^{1/2}$ with $\xi=0.005$.
We minimize the $\epsilon$-pseudospectral abscissa $\alpha_{\epsilon}(P(\cdot;\nu))$ for $(w_{m},w_{c},w_{k})=(1,1,1)$ and $\epsilon=0.05$ over $\nu\in[0,100]$ using EigOpt and Algorithm 1 to compute the objective $\alpha_{\epsilon}(\cdot;\nu)$ and its derivatives at various $\nu$.
Computing the globally optimal value with an error less than $10^{-8}$ requires 414 $\epsilon$-pseudospectral abscissa evaluations, and takes about 0.7 seconds. The computed global minimizer $\nu_{\ast}=4.6679$ (i.e., the optimal viscosity) is illustrated in the plots of $\alpha_{\epsilon}(P(\cdot;\nu))$ as a function of the viscosity $\nu$ on the left in Figure 2.
The original undamped system is asymptotically stable with the spectral abscissa $\alpha(P(\,\cdot\,;0))=-0.0043$, but not robustly stable as $\alpha_{\epsilon}(P(\,\cdot\,;0))=0.0619$ is positive indicating the existence of nearby systems that are not asymptotically stable. On the other hand, the globally minimal value of the $\epsilon$-pseudospectral abscissa $\alpha_{\epsilon}(P(\,\cdot\,;\nu_{\ast}))=-0.0888$ is negative, so not only this damped system but also all nearby systems are asymptotically stable. Moreover, the spectral abscissa for the optimized system $\alpha(P(\,\cdot\,;\nu_{\ast}))=-0.1347$ is smaller than that for the undamped system by a magnitude of more than thirty times, indicating also a better asymptotic behavior for the optimally damped system. The $\epsilon$-pseudospectrum of the original undamped system and the optimally damped system are shown on the right in Figure 2. While some of the components of the $\epsilon$-pseudospectrum for the undamped system extend to the right-half of the complex plane, all components of the $\epsilon$-pseudospectrum of the optimally damped system lie in the left-half of the complex plane. Indeed, optimization moves all of the components of the $\epsilon$-pseudospectrum considerably to the left.
Refer to caption Refer to caption
Figure 2: Small-scale damping optimization problem with $\epsilon=0.05$ in Example 5.1. The left plots illustrate $\alpha_{\epsilon}(P(\cdot;\nu))$ as a function of the viscosity $\nu$, as well as its computed global minimizer marked with red circles. In the bottom plot, the $\epsilon$-pseudospectral abscissa objective is zoomed near the computed global minimizer. In the right plot, the boundary of the $\epsilon$-pseudospectrum for the originally undamped system $\partial\Lambda_{\epsilon}(P(\cdot;0))$ (in blue) and that for the optimally damped system $\partial\Lambda_{\epsilon}(P(\cdot;\nu_{\ast}))$ (in green) are shown along with their eigenvalues.
Example 5.2.
A larger example similar to Example 5.1 involves $20\times 20$ mass, damping, stiffness matrices given by
$M=\mathrm{diag}(1,\dots,20)\>,\;\;C(\nu)=C_{\mathrm{int}}\,+\,\nu e_{2}e_{2}^{T}\>,\;\;K=\mathrm{tridiag}(-25,50,-25)\,,$
[…]
Refer to caption Refer to caption
Figure 3: Results for $20\times 20$ damping problem in Example 5.2.
The iterates $\nu^{(k)}$, $\alpha_{\epsilon}(\nu^{(k)})$ for $\,k\geq 1$ of Algorithm 3 listed in Table 2 indicate a fast convergence. The total runtime of Algorithm 3 until termination is about 1.8 seconds. This is much faster than a direct minimization (i.e., without employing subspace frameworks) of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with Algorithm 1 to compute the objective, which takes about 87.5 seconds. Even a direct minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$ via EigOpt coupled with the subspace framework (i.e. Algorithm 2) to compute the $\epsilon$-pseudospectral abscissa takes about 25.4 seconds and is still substantially slow. The distribution of the runtime among the different parts of Algorithm 3 is presented in the second, third, fourth columns of Table 5.2. Most of the runtime of Algorithm 3 is taken by the solutions of the reduced problems in line 3, and the computation of the rightmost point in $\Lambda_{\epsilon}(P(\cdot;\nu))$ via Algorithm 2 in lines 1 and 4. Among these two ingredients, solving the reduced problems contributes more to the total runtime of Algorithm 3.
$k$ $\nu^{(k)}$ $|\nu^{(k)}-\nu_{\ast}|$ $\alpha_{\epsilon}(\nu^{(k)})$
1 29.15901  $\,3.29\cdot 10^{1}$0.20179812
2 42.10685    $\;\;\;7.66\cdot 10^{-3}$0.00198921
3 42.10761 $\><10^{-5}$0.00199163
Table 2: Iterates of Algorithm 3 to solve the damping optimization problem in Example 5.2. The underlined digits indicate correct decimal digits.
  iter   total reduced a psa a inner direct (Alg. 1) direct (Alg. 2)
4 1.8 1.6 0.2 4, 6 87.5 25.4
Table 3: Number of iterations and runtimes in seconds to solve the damping problem in Example 5.2. The total time (the column ‘total’), time taken by the reduced problems inside Algorithm 3 (the column ‘reduced’), and time to compute $\alpha_{\epsilon}(P(\cdot;\nu))$, corresponding rightmost point inside Algorithm 3 (the column ‘psa’) are also listed. The column of ‘iter’ corresponds to the number of (outer) iterations of Algorithm 3, while, in the column of ‘inner’, the average number (rounded to nearest integer) and maximal number of (inner) iterations performed by Algorithm 2 during several calls throughout the minimization process via Algorithm 3 are given. The last two columns are times for direct minimization coupled with Algorithm 1 or Algorithm 2 to compute $\alpha_{\epsilon}(P(\cdot;\nu))$.
The inner subspace framework Algorithm 2 is called several times during this minimization via Algorithm 3. The average number of iterations (rounded to nearest integer) and maximal number of iterations of Algorithm 2 during these calls are four and six, respectively. We illustrate a typical convergence behavior of the inner subspace framework in Table 4, which lists the iterates of Algorithm 2 to compute $\alpha_{\epsilon}(P(\cdot;\nu_{\ast}))$. A superlinear convergence of the iterates is apparent from the table. Here, the inner subspace framework terminates after four iterations with a restriction subspace of dimension four. This is a typical convergence behavior which seems to be independent of the size of the matrix polynomial; the inner iteration terminates with a subspace of dimension less than ten in all of our experiments. See in particular the number of inner iterations reported below for the subsequent two examples.
$k$ $z^{(k)}$
[…]
Table 4: The iterates of Algorithm 2 on the damping problem in Example 5.2 to compute $\alpha_{\epsilon}(P(\cdot;\nu_{\ast}))$ at the global minimizer $\nu_{\ast}$.
Example 5.3.
To illustrate the behavior of the framework with increasing dimension $n$, this example concerns damping optimization problems with $n\times n$ mass, damping, stiffness matrices of the form
[…]
The computed results of Algorithm 3 are reported in Table 6. In each case, the $\epsilon$-pseudospectral abscissa of the optimally damped system is considerably smaller than that of the original undamped system. In some cases, even the $\epsilon$-pseudospectral abscissa becomes negative after minimization, indicating not only the optimal system but also nearby systems are made asymptotically stable through optimization. The number of subspace iterations of and runtimes of Algorithm 3 are also listed in Table 6. An important observation is that the number of subspace iterations of Algorithm 3 until termination remains small, indeed never exceeds three, in all cases. Similarly, according to the figures in the last column of Table 6, the average and maximal number of inner iterations of Algorithm 2 throughout an execution of Algorithm 3 never exceeds five and eight, respectively. The dimension of the restriction subspace for the outer subspace framework (Algorithm 3) is bounded by $\eta+\text{iter}-1$, which, recalling that $\eta=5$ initial interpolation points are used, is less than eight in all cases. Also, the dimension of the restriction subspace for the inner subspace framework (Algorithm 2) never exceeds eight according to the maximal number of inner iterations listed in the last column of Table 6.
[…]
Both the outer and the inner subspace frameworks are key ingredients for an efficient minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$. To illustrate this point, for the example with $n=80$, we report the total runtimes in Table 5 when both of the frameworks are employed, as opposed to when only one of the two subspace frameworks is adopted, while the other one is abandoned. The column ‘Alg. 3 (with Alg. 2)’ in the table is the approach proposed here when both of the subspace frameworks are employed, while the two columns to the right are when the inner framework Algorithm 2 is omitted, and when the outer framework Algorithm 3 is omitted. Not using either one of the frameworks increases the runtime considerably, but the effect of excluding the outer framework (i.e. Algorithm 3) on the runtime is more pronounced.
$w$ Alg. 3 (with Alg. 2) Alg. 3 (with Alg. 1) direct (with Alg. 2)
$\mathbf{1}$ 12.3 41.4 258.4
$\mathbf{u}$ 18.4 45.9 251.5
[…]
$n$, $\>w$ $\nu_{\ast}$ $\alpha_{\epsilon}(\nu_{\ast})$ $\alpha_{\epsilon}(0)$ iter total reduced psa inner
$80$, $\mathbf{1}$ 122.48 -0.00223 0.25226 3 12.3 11.7 0.5 3, 6
$80$, $\mathbf{u}$ 226.67 -0.00037 0.13030 3 18.4 17.9 0.5 3, 3
$200$, $\mathbf{1}$ 123.50 -0.00438 0.25226 3 73.2 67.9 5.0 4, 7
$200$, $\mathbf{u}$ 227.20 -0.00002 0.13030 3 111.4 107.1 4.2 3, 3
$400$, $\mathbf{1}$ 124.56 -0.00661 0.25226 3 260.1 227.1 33 5, 8
$400$, $\mathbf{u}$ 227.25 -0.00001 0.13030 2 142.7 120.1 22.4 3, 3
$1200$, $\mathbf{1}$ 125.48 -0.00850 0.25226 2 820.3 98.3 692.8 5, 7
$1200$, $\mathbf{u}$ 227.25 -0.00001 0.13030 2 774.6 134.9 618.5 3, 3
$1400$, $\mathbf{1}$ 125.48 -0.00850 0.25226 2 1209.1 98.2 1091.1 5, 6
$1400$, $\mathbf{u}$ 227.25 -0.00001 0.13030 2 1136.8 137.4 989.6 3, 3
Table 6: Application of Algorithm 3 to solve the damping problems of size $n=80,200,400,1200,1400$ in Example 5.3, and with the weights $w=(w_{m},w_{c},w_{k})$ equal to $\mathbf{1}:=(1,1,1)$ or $\mathbf{u}:=(0.7,1,0)$. The second, and third columns list the computed global minimizer, and the corresponding minimal value of $\alpha_{\epsilon}(\nu)$. The last five columns are as in Table 3.
Example 5.4.
We finally discuss two damping optimization problems that depend on two parameters. Both of the problems are variations of Example 5.2. In the first problem, the mass, damping, stiffness matrices are $20\times 20$ and given by
$M=\mathrm{diag}(1,\dots,20),\;C(\nu_{1},\nu_{2})=C_{\mathrm{int}}+\nu_{1}e_{2}e_{2}^{T}+\nu_{2}e_{19}e_{19}^{T},\;K=\mathrm{tridiag}(-25,50,-25),$
where $C_{\mathrm{int}}$ is as in Example 5.2. Thus, there are now two dampers on the second and nineteenth mass instead of only one damper on the second mass in Example 5.2. The optimization parameters $\nu_{1}$ and $\nu_{2}$, the viscosities of the dampers, are constrained to lie in the intervals $[0,50]$ and $[0,100]$, respectively, i.e. $\nu=(\nu_{1},\nu_{2})\>\in\>\underline{\Omega}=[0,50]\times[0,100]$. An application of Algorithm 3 to minimize $\alpha_{\epsilon}(\nu)$ for $\epsilon=0.05$ and $(w_{m},w_{c},w_{k})=(1,1,1)$ over $\nu\in\underline{\Omega}\,$ yields $\nu_{\ast}=(27.5958,62.1559)$ as the optimal values of the parameters. The accuracy of this computed minimizer is confirmed by the left-hand plot in Figure 4. The quantities related to this application of Algorithm 3 are summarized in Table 7. The computed globally minimal value of the $\epsilon$-pseudospectral abscissa is $\alpha_{\epsilon}(\nu_{\ast})=-0.0186$. Recall from the discussions following Example 5.2 that the $\epsilon$-pseudospectral abscissa of the original undamped system is $0.1324$, while with only one damper on the second mass the smallest $\epsilon$-pseudospectral abscissa possible is $0.0020$. Using a second damper moves all of the components of the $\epsilon$-pseudospectral abscissa completely to the open left-half of the complex plane, i.e. all of the systems at a distance of $\epsilon=0.05$ or closer are asymptotically stable for the optimal choices for the two dampers. This robust asymptotic stability is not attainable with only one damper on the second mass as illustrated in Example 5.2.
Refer to caption Refer to caption
Figure 4: Contour diagrams of the objective function $\nu=(\nu_{1},\nu_{2})\mapsto\alpha_{\epsilon}(\nu)$ for the two damping optimization problems in Example 5.4 depending on two parameters. In each plot, the red dot marks the computed global minimizer. Left and right plots correspond to first and second damping problems, respectively, considered in Example 5.4.
$\nu_{\ast}$ $\alpha_{\epsilon}(\nu_{\ast})$   iter   total reduced   psa inner
(27.60, 62.16) -0.01865 5 54.7 54.3 0.4 4, 6
Table 7: Damping optimization problem with two dampers in Example 5.4. The computed global minimizers, globally smallest value of $\alpha_{\epsilon}(\nu)$ by Algorithm 3, as well as the number of outer and inner subspace iterations, runtimes of Algorithm 3 in seconds are listed as in Table 6.
In the second problem, the mass, damping, stiffness matrices are again $20\times 20$ but now given by
$M=\mathrm{diag}(1,\dots,20)\>,\;\,C(\nu_{1},\nu_{2})=C_{\mathrm{int}}(\nu_{2})\,+\,\nu_{1}e_{2}e_{2}^{T}\>,\;\,K(\nu_{2})=\nu_{2}\mathrm{tridiag}(-1,2,-1)$
[…]
$\nu_{\ast}$ $\alpha_{\epsilon}(\nu_{\ast})$   iter   total reduced   psa inner
(72.46, 80) -0.00805 3 15.1 14.8 0.3 4, 6
Table 8: This table is similar to Table 7 but for the application of Algorithm 3 to the second damping problem in Example 5.4 with the damping viscosity and the spring constant as parameters.
6 Minimizing the pseudospectral abscissa of linearizations
In this section we discuss the minimization of the pseudospectral abscissa for the case when the quadratic eigenvalue problem is first linearized. Let us assume for this that the leading coefficient $M$ of the quadratic polynomial $P(\lambda)=\lambda^{2}M+\lambda C+K$ is invertible. This is often the case in real applications such as the damping optimization and brake squealing problems; see, e.g. [10, 34]. Then $P(\lambda)v=0$ holds if and only if
(28) ${\mathcal{A}}\left[\begin{array}[]{c}\lambda v\\ v\end{array}\right]=\lambda\left[\begin{array}[]{c}\lambda v\\ v\end{array}\right]\>,\;\>\text{where}\;{\mathcal{A}}:=\left[\begin{array}[]{cc}-M^{-1}&0\\ 0&I\end{array}\right]\left[\begin{array}[]{cc}C&K\\ I&0\end{array}\right]\in{\mathbb{C}}^{2\mathsf{n}\times 2\mathsf{n}}.$
This corresponds to the companion linearization of the polynomial $P(\lambda)$. There are many alternative linearizations considered in the literature (see, e.g. [12, 18, 19]). We adopt the one above, but the ideas here are applicable to other linearizations as well.
Clearly, the polynomial $P$ and the matrix ${\mathcal{A}}$ have the same set of eigenvalues, and the relation between their eigenvectors is apparent from (28). Hence, it seems plausible to consider the $\epsilon$-pseudospectrum of the matrix ${\mathcal{A}}$ given by [35]
$\Lambda^{\mathsf{st}}_{\epsilon}({\mathcal{A}})\>:=\>\;\bigcup_{\Delta\in{\mathbb{C}}^{2\mathsf{n}\times 2\mathsf{n}}\,,\;\|\Delta\|_{2}\leq\epsilon}\,\Lambda({\mathcal{A}}+\Delta)\\ \quad=\quad\{z\in{\mathbb{C}}\;|\;\sigma_{\min}({\mathcal{A}}-zI)\leq\epsilon\}\>$
and the corresponding $\epsilon$-pseudospectral abscissa
$\begin{split}\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}})\>&:=\>\max\{\text{Re}(z)\;|\;z\in\Lambda^{\mathsf{st}}_{\epsilon}({\mathcal{A}})\}\\[1.00006pt] &\phantom{:}=\;\max\{\text{Re}(z)\;|\;z\in{\mathbb{C}}\text{ s.t. }\sigma_{\min}({\mathcal{A}}-zI)\leq\epsilon\}\>.\end{split}$
The next theorem relates $\Lambda_{\epsilon}(P)$ with $\Lambda^{\mathsf{st}}_{\epsilon}({\mathcal{A}})$, as well as $\alpha_{\epsilon}(P)$ with $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}})$.
Theorem 6.1.
Consider the quadratic matrix polynomial $P(\lambda)$ with $M$ invertible, and let $\Lambda_{\epsilon}(P)$, $\alpha_{\epsilon}(P)$ be the $\epsilon$-pseudospectrum, $\epsilon$-pseudospectral abscissa of $P$ with the weights $(w_{m},w_{c},w_{k})=(0,1,1)$. The following assertions hold for every $\epsilon>0$:
1. $\Lambda^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}})\supseteq\Lambda_{\epsilon}(P)$;  2. $\alpha^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}})\geq\alpha_{\epsilon}(P)$,  where $\underline{\epsilon}:=\epsilon\|M^{-1}\|_{2}$.
Proof 6.2.
1. Suppose that $z\in\Lambda_{\epsilon}(P)$. Then there must exist $\Delta C,\Delta K$ such that $\|[\Delta C\;\Delta K]\|_{2}\leq\epsilon$ and $z$ is an eigenvalue $P(\lambda)+\Delta P(\lambda)$ for $\Delta P(\lambda)=\lambda\Delta C+\Delta K$. But then $z$ is also an eigenvalue of the matrix
$\begin{split}&\left[\begin{array}[]{cc}-M^{-1}&0\\ 0&I\end{array}\right]\left[\begin{array}[]{cc}C+\Delta C&K+\Delta K\\ I&0\end{array}\right]={\mathcal{A}}+\Delta\\ &\quad\quad\text{for}\quad\Delta:=\left[\begin{array}[]{cc}-M^{-1}&0\\ 0&I\end{array}\right]\left[\begin{array}[]{cc}\Delta C&\Delta K\\ 0&0\end{array}\right]\>,\end{split}$
where $\|\Delta\|_{2}\;\leq\;\|M^{-1}\|_{2}\|[\Delta C\;\Delta K]\|_{2}\;\leq\;\epsilon\|M^{-1}\|_{2}\>.$ This shows that $z\in\Lambda^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}})$ for $\underline{\epsilon}:=\epsilon\|M^{-1}\|_{2}$. Hence, the inclusion $\Lambda^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}})\supseteq\Lambda_{\epsilon}(P)$ holds.
2. This is immediate from the inclusion $\Lambda^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}})\supseteq\Lambda_{\epsilon}(P)$.
Returning to the parameter-dependent matrix polynomial $P(\lambda;\nu)=\lambda^{2}M(\nu)+\lambda C(\nu)+K(\nu)$ under the assumption that $M(\nu)$ remains invertible for all $\nu\in\underline{\Omega}$, letting
${\mathcal{A}}(\nu):=\left[\begin{array}[]{cc}-M(\nu)^{-1}&0\\ 0&I\end{array}\right]\left[\begin{array}[]{cc}C(\nu)&K(\nu)\\ I&0\end{array}\right]\>,$
one may alternatively minimize $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$ over $\nu\in\underline{\Omega}$ based on the stability considerations. This involves the minimization of the $\epsilon$-pseudospectral abscissa of the parameter-dependent matrix ${\mathcal{A}}(\nu)$, so falls into the setting of [2, 24].
If the weights in the definitions of $\Lambda_{\epsilon}(P(\cdot;\nu))$ and $\alpha_{\epsilon}(P(\cdot;\nu))$ are chosen as $(w_{m},w_{c},w_{k})=(0,1,1)$, and $M(\nu)=M$ is independent of the parameters $\nu$, then, according to Theorem 6.1, we have
$\min_{\nu\in\underline{\Omega}}\>\alpha_{\epsilon}(P(\cdot;\nu))\;\;\leq\;\;\min_{\nu\in\underline{\Omega}}\>\alpha^{\mathsf{st}}_{\underline{\epsilon}}({\mathcal{A}}(\nu))$
with $\underline{\epsilon}=\epsilon\|M^{-1}\|_{2}$. However, the gap between the optimal values of the two minimization problems above can be significant, and the corresponding optimal parameter values can be considerably different; see in particular Example 6.3 given next.
Example 6.3.
For a comparison of $\alpha_{\epsilon}(P(\cdot;\nu))$ and $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$, consider the $20\times 20$ damping optimization problem in Example 5.2 with $\epsilon=0.05$ and the interval $\underline{\Omega}=[0,100]$. In the definition of $\alpha_{\epsilon}(P(\cdot;\nu))$, the weights are chosen as $(w_{m},w_{c},w_{k})=(0,1,1)$. Moreover, as $\|M^{-1}\|_{2}=1$, we have $\underline{\epsilon}=\epsilon\|M^{-1}\|_{2}=\epsilon$, so it follows from Theorem 6.1 for that all $\nu\in\underline{\Omega}$ we have $\alpha_{\epsilon}(P(\cdot;\nu))\leq\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$.
Minimizing $\alpha_{\epsilon}(P(\cdot;\nu))$ over $\nu\in\underline{\Omega}$ using Algorithm 3 yields $\nu_{\ast}=66.42$ as its global minimizer, and $\alpha_{\epsilon}(P(\cdot;\nu_{\ast}))=0.0012$ as the globally smallest value of $\alpha_{\epsilon}(P(\cdot;\nu))$. On the other hand, the minimization of $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$ over $\nu\in\underline{\Omega}$ using the framework in [2], in particular the software in [24], leads to $\nu^{\mathsf{st}}_{\ast}=55.47$ as the global minimizer, and $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu^{\mathsf{st}}_{\ast}))=0.1155$ as the globally smallest value of $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$.
The accuracy of these computed results is illustrated by the plots of $\alpha_{\epsilon}(P(\cdot;\nu))$ and $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$ in Figure 5. The computed global minimizers $\nu_{\ast}=66.42$ and $\nu^{\mathsf{st}}_{\ast}=55.47$ are away from each other. Moreover, $\alpha^{\mathsf{st}}_{\epsilon}({\mathcal{A}}(\nu))$ is considerably larger than $\alpha_{\epsilon}(P(\cdot;\nu))$ by a magnitude of at least ten globally for all $\nu\in\underline{\Omega}$. A comparison of the spectral abscissa $\alpha(\nu_{\ast})=-0.00844$ and $\alpha(\nu^{\mathsf{st}}_{\ast})=-0.00843$ at the computed global minimizers indicates that the system $M\ddot{x}(t)+C(\nu)\dot{x}(t)+Kx(t)=0$ has slightly better asymptotic behavior at $\nu=\nu_{\ast}$ compared to $\nu=\nu^{\mathsf{st}}_{\ast}$.
Refer to caption Refer to caption
Figure 5: Comparison of $\alpha_{\epsilon}(P(\cdot;\nu))$ with $\alpha_{\epsilon}^{\mathsf{st}}({\mathcal{A}}(\nu))$ on the damping optimization problem in Example 5.2. In both plots, the red circle marks the computed global minimizer. Left Plot: $\alpha_{\epsilon}(P(\cdot;\nu))$ as a function of $\nu$ for $(w_{m},w_{c},w_{k})=(0,1,1)$. Right Plot: $\alpha_{\epsilon}^{\mathsf{st}}({\mathcal{A}}(\nu))$ as a function of $\nu$.
7 Concluding Remarks
We have proposed numerical approaches for the minimization of the $\epsilon$-pseudospecrtral abscissa $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ over the parameters $\nu$ that belong to a compact set $\underline{\Omega}\in{\mathbb{R}}^{\mathsf{d}}$ for a given parameter-dependent quadratic matrix polynomial $P(\,\cdot\,;\nu)$ and for a prescribed $\epsilon>0$. When $P(\,\cdot\,;\nu)$ is of small size, the minimization of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ is performed by means of a global optimization technique that can cope with the nonsmoothness of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$ (or alternatively with BFGS if there are many optimization parameters) coupled with a globally convergent criss-cross algorithm to compute the objective $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$. Otherwise, when $P(\,\cdot\,;\nu)$ is not small, it is minimized by applying a subspace framework that operates on a restriction of $P(\,\cdot\,;\nu)$ to a small subspace ${\mathcal{V}}$. We have provided formal arguments that indicate the convergence of this framework to a global minimizer of $\alpha_{\epsilon}(P(\,\cdot\,;\nu))$. Numerical experiments on damping optimization examples show that the proposed methods can cope with the minimization of $\alpha_{\epsilon}(P(\cdot;\nu))$ accurately and efficiently for problems of size at least a few hundreds in practice.
Efficient and reliable $\epsilon$-pseudospectral abscissa computations in the large-scale setting are one of the main obstacles standing in the way to deal with even larger $\epsilon$-pseudospectral abscissa minimization problems for quadratic eigenvalue problems, e.g. those related to the robust stability of finite element discretizations of a disk brake [10]. The framework we rely on (i.e. Algorithm 2) for these larger $\epsilon$-pseudospectral abscissa computations is a remedy, but still needs improvements to avoid local convergence such as a clever initialization strategy, possibly making use of the perturbation theory for quadratic eigenvalue problems.
Acknowledgements. The authors are grateful to two anonymous referees for their comments and suggestions, which helped the authors to improve the manuscript.